### v3

# Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation

Abstract
Recently, Large Language Models (LLMs) have shown impressive performance in character understanding tasks, such as analyzing the roles, personalities, and relationships of fictional characters. However, the extensive pre-training corpora used by LLMs raise concerns that they may rely on memorizing popular fictional works rather than genuinely understanding and reasoning about them. In this work, we argue that “gist memory”—capturing essential meaning—should be the primary mechanism for character understanding tasks, as opposed to “verbatim memory”—exact match of a string. We introduce a simple yet effective method to mitigate mechanized memorization in character understanding evaluations while preserving the essential implicit cues needed for comprehension and reasoning. Our approach reduces memorization-driven performance on popular fictional works from 96% accuracy to 72% and results in up to an 18% drop in accuracy across various character understanding tasks. These findings underscore the issue of data contamination in existing benchmarks, which often measure memorization rather than true character understanding.
Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation
Yuxuan Jiang , Francis Ferraro Department of Computer Science and Electrical Engineering University of Maryland, Baltimore County yuxuanj1@umbc.edu
1 Introduction
Character understanding requires the models to grasp the characters’ personalities and traits while reading a story, which aims to evaluate the models’ capability to comprehend personalities, goals, roles, and relationships of various fictional characters Yu et al. (2023a); Vallurupalli et al. (2024); Shen et al. (2024); Brahman et al. (2021). However, the impressive performance of LLMs on such benchmarks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023) raises the question: does it reflect genuine understanding or merely strong memorization Chang et al. (2023a)? For example, in the TV show speaker guessing task shown in Figure  1, the model is expected to identify Sheldon based on its memory of the character from previous episodes (e.g., preparing lessons for Howard). However, if the model simply memorizes the exact words, it could make predictions based on a very short sentence by directly extracting the speaker information from it.
To truly stimulate a model’s character understanding ability, it is crucial to understand how the model may be generating its responses. This can be especially challenging given the extensive and opaque pre-training corpora used by LLMs, making it particularly severe for character understanding tasks, as many benchmarks are derived from well-known and widely consumed fiction or scripts Sang et al. (2022a). We hypothesize that this can lead to a form of memorization called “verbatim memory,” which encodes precise details and remembers every word exactly Huang et al. (2024); Ippolito et al. (2022)—rather than reasoning over recalled overarching patterns (what is referred to as “gist memory” Reyna and Brainerd (1998)).
Figure 1: This figure illustrates the distinct mechanisms of gist and verbatim memory processes in the speaker-guessing task for a script. Verbatim memory (bottom right) aims to recall as much specific information as possible, including who said each sentence. If the memorization is not precise enough, the model fails to provide the correct answer. In contrast, gist memory (top right) focuses on extracting core information, such as character relationships, personalities, and key events. By relying on gist memory, the model is encouraged to engage in reasoning rather than purely relying on memorization.
Motivated by these concepts, we propose an approach that alters elements such as character names, place names, and specific notations in the sentences while preserving the relationships between characters, significant events, and other information we want the model to utilize. We hypothesize that these modifications can reduce the model’s tendency to take shortcuts by relying on verbatim memory, and instead encourages the model to rely on gist memory to stimulate its reasoning ability for character understanding tasks. We tested our method on six key types of character understanding tasks and obtained an interesting finding: When the difficulty of using verbatim memory increases, the model is more likely to resort to reasoning to compensate.
Our work provides a simple approach to ensure a fairer benchmark for character understanding tasks on large models that already have strong memorization abilities of fictional works. It offers an alternative method compared to more costly solutions, such as developing data from obscure or very new books that the model has not encountered Ngo et al. (2024), or using techniques like machine unlearning to forget original training texts Shaik et al. (2024).
In summary, we provide an outline of the paper and summarize our key contributions:
  • To determine whether the memory types used in character understanding tasks align with our expectations, we draw on human cognitive science to define gist and verbatim memory in the context of character understanding tasks (s 2).
  • We benchmarked the mechanism memorization ability of over 100 sources of fictional works (s 4.1).
  • We measured the verbatim memory of large models for hundreds of different source fiction works and tested how different keywords affect the verbatim memory of these models (s 4).
  • We demonstrated that even for the same factor, such as a person’s name, different disturbances, such as gender or cultural interference, can lead to significantly different verbatim memory effects in large models (s 5).
  • We compared the different effects of the model using verbatim and gist memory across six different test sets and found that the model’s accuracy decreased by 0.7% to 17.2% (s 6).
2 Related Task
Among Large Language Models (LLMs)’ various capabilities Li et al. (2024b, d); Zhang et al. (2024); Li et al. (2024c), the reasoning abilities Tong et al. (2023); Yang et al. (2023); Xiong et al. (2024b); Tong et al. (2024); Xiong et al. (2024a) of LLMs Li et al. (2024a); Tan et al. (2024) have garnered significant attention, but their memory has also been shown to play a crucial role in influencing benchmark reasoning performance Zhou et al. (2023). This influence is complex: data contamination can inflate performance metrics Khandelwal et al. (2020); Blevins and Zettlemoyer (2022); Carlini et al. (2023), while knowledge conflict can reduce them Su et al. (2024); Wang et al. (2024); Xie et al. (2024b); Xu et al. (2024). Recent efforts have aimed to mitigate the impact of memory on reasoning benchmarks to establish more reliable evaluations, particularly for reasoning tasks like numerical reasoning, commonsense reasoning, or logical reasoning, where memory requirements are less significant and easier to define Xie et al. (2024a); Sainz et al. (2023); Li et al. (2024e). However, in character understanding tasks, the diversity of task formats and scenarios means that memory requirements vary significantly, making it challenging to establish a universal standard to distinguish between acceptable and unacceptable use of memory. We propose leveraging gist memory to reduce LLMs’ reliance on precise memorization. This approach encourages models to utilize reasoning capabilities more effectively in solving character understanding tasks, aligning them more closely with the original intent behind these benchmarks.
Character Understanding
Character understanding has made significant contributions to advancing story comprehension in psychological, literary, and educational research  Bower and Morrow (1990); McKee (2010); Paris and Paris (2003). This multifaceted field encompasses a variety of tasks, including character recognition, coreference resolution, summarization, role detection, goal understanding, personality analysis, and question answering Sang et al. (2022b); Bamman et al. (2020); Chen et al. (2022); Frermann et al. (2018); Vallurupalli et al. (2024); Yu et al. (2023b). Recent studies have shown promising performance by large language models (LLMs) on character understanding tasks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023). However, these studies often overlook the potential influence of memorization on model performance. Fictional works, including movie and TV show scripts, famous novels, and news texts, serve as primary sources for narrative understanding Sang et al. (2022a); Tapaswi et al. (2016); Caselli and Vossen (2017). These texts, much like copyrighted books Chang et al. (2023a), are prone to strong memorization. In our work, we measured the verbatim capabilities of LLMs on these sources, as well as the potential inflation of benchmark results caused by overly strong verbatim memory. Our approach also mitigates this effect.
Verbatim and Gist Memory
Cognitive science defines two types of memory: verbatim memory and gist memory Brainerd and Reyna (2002); Reyna and Brainerd (1998). Verbatim memory recalls the exact wording of a sentence, such as "The black-and-white cat sat on a soft, red mat in the living room," retaining every detail about the cat’s color, the mat’s texture and color, and the location. In contrast, gist memory captures the essential meaning, like "A cat was sitting indoors," focusing on the core idea—the presence of a cat and its general location—while discarding details such as the cat’s color or the specific description of the mat. This highlights how verbatim memory retains all specific details, while gist memory abstracts the key information by ignoring less critical details. Cognitive science research suggests that humans primarily rely on gist memory for reasoning, and these two types of memory are independent, rather than hierarchical Brainerd and Reyna (1992). In recent studies on LLM memorization Carlini et al. (2023), the extraction of strings from a larger text is defined as a form of verbatim memory. Their solution primarily focuses on generating entirely new test samples that the model has never encountered Xie et al. (2024a). However, for character understanding tasks, using obscure or newly released datasets that the model hasn’t seen comes with high costs, and the quality and acquisition costs of the texts are difficult to control. Our work demonstrates that using gist memory provides a simple and effective way to alleviate this issue.
3 Data and Experiment Setup
In our experiments, we utilized two distinct datasets: (1) Benchmark Datasets: six pre-existing datasets designed to evaluate various character understanding tasks, and (2) Fictional Works Dataset: a newly constructed dataset consisting of script excerpts sampled directly from the original texts used to create the benchmark datasets. This dataset, drawn from 40 diverse sources—including TV shows, novels, and movies—was specifically designed to investigate GPT’s familiarity with the original texts underlying the benchmark datasets.
3.1 Benchmark Datasets
We evaluated the interplay between memory and reasoning across six types of character understanding tasks. The datasets used are detailed in this section, while task descriptions are provided in Section 6.1.
Datasets
For coreference resolution , we use the latest released version of the MovieCoref dataset  [^1] https://github.com/usc-sail/mica-character-coref. For character guessing, we adopt the TVSHOWGUESS dataset  [^2] https://github.com/YisiSang/TVSHOWGUESS to conduct experiments. For personality understanding, we utilize the PERSONET dataset  [^3] https://github.com/Gorov/personet_acl23. For role detection, we use the CSI dataset  [^4] https://github.com/EdinburghNLP/csi-corpus. For open-domain question answering, we conduct experiments using the FriendsQA dataset  [^5] https://github.com/emorynlp/FriendsQA. For summarization task, we adopt the SummScreen dataset  [^6] https://github.com/mingdachen/SummScreen. We follow all the data preprocessing, development, and testing separation provided by the original datasets. Tasks and datasets details are in appendix G.
Evaluation Metrics
We followed the original evaluation metrics used in the respective tasks. For classification tasks (Character Guessing, Question Answering and Personality Understanding), we report accuracy using exact match. For sequence labeling and coreference tasks (Coreference and Role Detection), we use F1 scores as the evaluation metric. For summarization tasks (Text Summarization), we report ROUGE scores to measure the overlap between the generated and reference summaries.
Task Name Task Type Data Source Metrics
TVshow Guessing Character Guessing TV Shows Exact Match
FriendsQA Question Answering TV Shows Exact Match
Screenplay Character Coreference Movies F1 Score
SummScreen Text Summarization TV Shows ROUGE
CSI Role Detection TV Shows F1 Score
PERSONET Personality Understanding Novels Exact Match
Table 1: Overview of benchmark tasks, including their type, data sources, and evaluation metrics.
3.2 Fictional Works Dataset
The existing six benchmark datasets, which are derived from 91 different TV shows, 12 different movies, and 33 classic literature books. We constructed the Fictional Works Dataset by uniformly sampling 200 segments from 40 selected sources, including 11 distinct TV series, 9 movies, and 20 novels, each source provides 5 scripts. Each segment contains an average of 312 words. The format of each segment is: TV Show Title; Scene, where the scene consists of a mixture of dialogue and narration from the script.
3.3 Extended Dataset Analysis
For the six representative character understanding tasks studied in our work, we selected one dataset for each category. However, we also recognize that there are many different benchmark datasets for each type of task. When selecting representative datasets for each category, we focused on choosing those with more diverse sources. Additionally, we supplemented the selection with eight datasets of the same type but derived from different sources in Table 2.
Task Name Task Type Source
Character Mining Character Identification TV Shows
LISCU Character Identification Book Reviews
MovieQA Visual Question Answering Movies
Liter Coreference Character Coreference Novels
HVV Detect Role Detection Movies, News
SYMON Visual Question Answering Movies
CEB Character Representation Novels
Love and Violence Summarization TV Shows
Table 2: Overview of 14 tasks in the character understanding benchmark.
3.4 Model
We pass all passages with the same prompt through GPT4o (gpt-4o), [^7] https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo using the OpenAI API. The cost for the tasks in s 4 and s 5 is approximately \$80, while the cost for the task in s 6 is approximately \$40. The character description used in section  5 is generated by GPT-4 (gpt-4-turbo) API. For the prompts and description generated by GPT4 itself, please refer to Appendix C. Due to the inherent characteristic of ChatGPT’s unstable output, even though we had set the temperature to 0, each generation may exhibit slight variations in content. In our experiments, there is a variation of 1% to 3% in the results. To mitigate this, we repeat each experiment three times and report the average scores.
4 Verbatim Memorization Evaluation
We address our first research question: How does GPT utilize and integrate multi-source information to identify the origin of a script passage?
We do this by considering a task that is focused on evaluating the extent to which models rely on verbatim memory to solve problems without requiring significant reasoning.
4.1 Task Definition
Name Source Prediction with Rationale
In this Source Memorization task, we assess the model’s ability to predict the origin of a text passage, investigating the extent to which it relies on memorized pretraining knowledge. The model is given a segment script and we ask it to identify the source while also providing a rationale for its prediction 9. This task includes two experimental settings: (1) with or without multiple-choice options, and (2) with or without the requirement for a rationale explanation, which made four combinations.
Rationale Analysis
We collected and analyzed the rationales generated by GPT, then asked GPT to performing clustering on all rationales and find groups of rationales with similar pattern. This analysis helps identify common elements the model relies on during source prediction, especially when character names are explicitly mentioned as part of the reasoning.
Effect of Input Length
To explore the impact of input length, we limited the length of text segments provided to the model. We compared its performance in two settings: one with character names and specific mentions provided, and the other with them masked. We hypothesize that verbatim memory plays a lesser role in such tasks, aligning with the human prediction process Brainerd and Reyna (1992).
4.2 Result
Source Prediction
The results presented in Table 3 reveal several notable trends across the different experimental setups. First, the model demonstrates a strong memorization ability when presented with the original script, achieving a consistent 96% accuracy even without any options and under the Chain-of-Thought (CoT) condition. This suggests a significant reliance on memorized content, as the model can accurately retrieve the source purely based on internal knowledge. Furthermore, the use of CoT reasoning substantially enhances the model’s performance across all conditions, improving the retrieval accuracy from 96% to 100% in the case of the original script. This indicates that CoT prompts help guide the model toward more accurate predictions by encouraging a more structured reasoning process. However, when character names are masked (using placeholders like P0, P1), the model’s accuracy decreases, highlighting the importance of specific names as primary retrieval cues. The drop in performance is even more pronounced when names are replaced with generic alternatives, suggesting that the model’s retrieval capability is highly dependent on the exact memorization of character names rather than broader contextual information. This is evidenced by the reduction in accuracy from 96% with the original script to 74% with replaced names. In contrast, replacing notable mentions (such as unique terms or identifiable elements other than character names) has a minimal effect on performance, with the model maintaining a high accuracy of 96%-98%. This further underscores that the model’s retrieval strategy is heavily focused on character names, and it places less emphasis on other narrative details when determining the text’s origin. Overall, these findings highlight the model’s dependency on character-specific memorization, revealing both the strengths and vulnerabilities of its retrieval mechanisms when faced with modifications to key identifying elements.
Condition With Options CoT No Options CoT
Origin Script 96% 100% 96% 98%
Name Masked (P0, P1) 92% 98% 82% 86%
Name Replaced 84% 98% 74% 80%
Notable Mentions Replaced 94% 98% 96% 96%
Name + Mention Replaced 84% 96% 72% 80%
Table 3: We find that CoT consistently improves the model’s prediction accuracy. Character names contribute significantly to accuracy, while notable mentions have a very small impact, likely due to the correlation between their influence and their frequency of occurrence in the text.
Rationale Statistics
We analyzed the 7 rationale content generated by the model across these conditions. The following table (Table 4) shows the frequency of the key rationale elements identified in the model’s output. The most frequently mentioned rationale elements include character name, content of dialogue, and setting, indicating that these factors play a significant role in the model’s reasoning process.
Rationale Element Percentage (%)
Character Name 84%
Notable Mentions 28%
Content of Dialogue 66%
Tone 40%
Humor Style 46%
Setting 46%
Social Dynamics 36%
Table 4: We found that the model primarily relies on the combinations of character names mentioned in the text, especially the main characters, when guessing the source of the text. Although notable mentions are rarely utilized, some distinctive catchphrases, special place names, and event titles are still noticed and leveraged by the model.
Word Limitation
Table 5 shows the model’s source identification accuracy with varying token limits, comparing the original script and the script with replaced names. The model consistently achieves 96% accuracy with the original script, demonstrating strong memorization even with limited context (as few as 25 tokens). In contrast, replacing character names leads to a significant drop in accuracy, especially with shorter passages. For the 25-token input, accuracy falls to 26%, highlighting the model’s heavy reliance on specific names for retrieval. Increasing the token limit to 50 tokens improves accuracy to 48%, but extending to 100 tokens shows no further gains. On average, across all 342 samples, the accuracy for replaced names recovers to 74%, suggesting some benefit from longer contexts but still falling short of the original script’s performance. This indicates that character names play a crucial role in the model’s retrieval process, with limited capacity to compensate through broader context when explicit names are obscured.
Word Limit Original Script Replaced Names
25 Words 96% 26%
50 Words 96% 48%
100 Words 96% 52%
All (Avg. 342) 96% 74%
Table 5: By leveraging verbatim memory, GPT demonstrates a surprisingly high memorization ability even with as few as 25 words. The slower improvement as the number of words increases suggests that source prediction predominantly relies on verbatim memory.
4.3 Analysis
Using text excerpts to prompt the model to guess their source highlights the model’s reliance on verbatim memory. Experiments with word count limitations revealed that the number of words had little impact on performance, further confirming that the model relies on verbatim memory for this task. By employing CoT, we encouraged the model to use gist memory for reasoning, resulting in improved performance and greater robustness against perturbations. This, in turn, provides additional evidence that the model predominantly relies on verbatim memory to solve such tasks.
5 Gist Memory Evaluation
The experiment in Section s 4 demonstrated that applying Chain-of-Thought (CoT) can prompt LLMs to utilize gist memory. In this section, we aim to explore additional methods to encourage LLMs to leverage gist memory effectively.We hypothesize that by reducing the confidence LLMs place in their verbatim memory, they are more likely to rely on gist memory while simultaneously applying reasoning. Building on previous findings that changing character names can significantly degrade performance, we intuitively believe that altering names allows the model to recognize changes in the information embedded within names, such as nationality and gender. Additionally, we pose our second question: Does name replacement affect LLMs’ reliance on gist memory, and if so, to what extent? Moreover, does the choice of name replacement method make a difference?
5.1 Task Definition
We adopt the character guessing  Sang et al. (2022b) format, and we masked the character names mentioned in the script as well. For the prompts please refer to Appendix D.
Name Replacement Strategies
We employed the following name replacement methods:
  • Cross-Language Replacement: Replacing names with uncommon names from other languages, swapping the gender in some cases (GS).
  • Same-Language Replacement: Replacing names with common names from the same language, including gender-swapped (GS) variants.
  • Anonymized Replacement: Replacing all character names with neutral labels (e.g., "Speaker 0").
  • Gender-Swap Only (GS Only): Swapping the gender of the original names without changing the language.
  • Control Tests: Introducing incorrect show names or names from other TV shows to evaluate the model’s susceptibility to data contamination (an example in  11).
5.2 Result
Our analysis focuses on understanding how different replacement strategies affect the model’s predictions. Gender-swapping and cross-cultural name replacements may disrupt the model’s memorized associations more effectively than simple anonymization. This helps us gauge the model’s inference capabilities when it cannot rely solely on pretraining knowledge. Table 6 summarizes the accuracy of GPT-4o across different replacement strategies with and without character descriptions. The "Name Return" task shows that GPT-4o achieves a perfect 100% accuracy, indicating strong memorization of original character names. However, the accuracy significantly drops for other replacement strategies, especially for cross-language and gender-swapped (GS) replacements, where the performance decreases to as low as 20.2% and 15.5%, respectively. Interestingly, anonymization (replacing all names with neutral labels like "Speaker 0" and "Speaker 1") shows a moderate increase in accuracy when descriptions are provided, suggesting the model may rely more heavily on contextual information in these cases. The results indicate that providing character descriptions can slightly improve the model’s performance, particularly in the cases of anonymized and incorrect show names replacements.
Replacement Strategy Without Description With Description
Name Return 100 100
No Replacement 52.1 47.9
Cross-Language 35.5 28.2
Cross-Language (GS) 20.2 21.8
Same Language 34.2 24.1
Same Language (GS) 15.5 16.5
Anonymized 16.7 26.4
GS Only 7.3 8.9
Wrong show names 41.5 43.6
Names from other shows 12.3 31.1
Table 6: The impact of different name replacement methods on GPT-4o’s verbatim memory ability was demonstrated. With reduced confidence in its verbatim memorization, GPT-4o is more likely to rely on reasoning. By providing descriptions, GPT-4o is more inclined to access gist memory and then apply reasoning.
5.3 Analysis
Providing character summarization evaluates two key aspects of GPT’s performance. Without summarization, the task primarily tests verbatim memory, where GPT tends to struggle. However, when character descriptions are provided, the task evaluates both memorization and inference abilities. We observe that as the task complexity increases, GPT relies more on reasoning and less on simple memorization.
6 Reasoning Validation Task
The character guessing task format is quite similar to verbatim memorization evaluation tasks, such as the name cloze task Chang et al. (2023b). We have demonstrated that tasks heavily reliant on verbatim memory can measure its reliance and help transfer the model’s usage to gist memory. Intuitively, most types of character understanding tasks should rely more on gist memory. This leads us to our third question: Can we encourage GPT to rely more on gist memory in different character understanding tasks?
6.1 Benchmark Tasks Description
Generally, character understanding tasks require the model to predict a character’s information based on a given segment of text. For script-based character understanding, the provided texts often include conversations from TV show and movie scripts, while novels typically include a mix of narrative descriptions and dialogue. The six selected benchmarks cover a range of character-centric tasks, designed to test different aspects of character understanding(Table 1):
Character Guessing This task involves identifying the speaker for each utterance in a script. Each utterance within a scene is segmented and input into the model, with the speaker’s name preceding each utterance masked and replaced by a unique special token. Consistent special tokens are used to represent the same speaker within a scene. The objective of the character guessing task is to accurately predict the identity of the speaker associated with each special token.
Coreference Resolution Given a conversation in scripts that contains multiple utterances and n character mention entity c1, c2, …, cn within it, the objective of the coreference resolution task is to assemble all mention entities that refer to the same character in a cluster.
Personality Understanding Given a text snippet describing a character from classic literature along with preceding context, the objective of the personality understanding task is to predict the most likely personality trait depicted in the snippet. The task is formulated as a multi-choice problem with distractor traits, testing the model’s ability to infer character personalities from descriptive text.
Role Detection Given dialogues and scene descriptions from scripts, the objective is to determine whether each token refers to the perpetrator. This task is framed as a multi-label classification problem, where each label corresponds to a character in the script, with a focus on detecting roles within a crime narrative context.
Open Domain Question Answering The input is a dialogue excerpt from the script and a manually crafted question, the objective of the open-domain question-answering task is to extract the specific span in the dialogue that answers the question. This can be viewed as an answer span selection task, where the model is expected to identify spans within the dialogue that contain the answer to the question.
Summarization Given dialogue utterances and scene descriptions from transcripts, the objective of the summarization task is to generate a coherent summary that captures key plot points while excluding irrelevant details.
6.2 Perturbed Method
We implemented the name-changing strategy of swapping genders across cultures, which, as demonstrated in Section s 5, significantly drops performance.
6.3 Result
Table 7 presents the impact of replacing character names on a variety of downstream tasks using the GPT-4o model. The experiments assess performance across six tasks:
Memory-Driven Tasks:
These tasks rely heavily on verbatim memory and are significantly affected by name replacement.
Intuively, TV Show Guessing and Screenplay Coreference should rely more on verbatim memory and they drop of 14.6% and 8.7% meets our expectation.
Reasoning-Driven Tasks:
These tasks primarily rely on gist memory and are more likely to engage in reasoning without external hints(CoT or extra description).
The slight drops in FriendsQA and Personet personality understanding tasks, with a decrease of 0.7% and 2%, align with our expectations. The 17.2% drop in role extraction is unexpectedly high, resembling the trend seen in coreference resolution tasks. The drop in ScreenSum is moderate, falling within expected ranges.
Figure 2: All tasks exhibited performance degradation. The decline in QA and personality understanding tasks was minimal, indicating that GPT effectively leveraged its reasoning abilities for these tasks. The most significant drop was observed in role detection, exceeding our expectations, which suggests that verbatim memory plays a substantial role in supporting this type of task.
Task Name Origin Script Name Replaced
TVShow Guessing 44.8 30.2
Screenplay Coreference 58.3 49.6
FriendsQA 29.3 28.6
ScreenSum (ROUGE_L) 19.1 13.8
CSI Role Detection 50.1 32.9
PERSONET Personality 40.0 38.0
Table 7: All tasks showed a decline in performance, with the smallest drop in QA, which exceeded our expectations. We believe this is because the questions asked were relatively simple and basic. There was almost no change in personality understanding, indicating that the model knows it should use gist memory for such tasks. The largest decline was observed in role extraction, which also exceeded our expectations, suggesting that the model heavily relies on verbatim memory when handling this type of task.
6.4 Analysis
Our experiment, through name replacement, increases the difficulty of the model relying on verbatim memory. As a result, the model is more likely to focus on gist memory, thereby enabling reasoning. By comparing the performance before and after name replacement, we can observe whether the model relies more on verbatim or gist memory without external hints. Specifically, if the performance difference is small, it suggests that the model is more likely to use reasoning for answering such tasks. Conversely, if the difference is large, it indicates that the model tends to rely more on verbatim memory. All tasks showed a decline, and we observed that the model always tends to prioritize verbatim memory. This ’lazy’ behavior suggests that to unlock the model’s true reasoning ability, we need to design methods that encourage the model to focus more on gist memory, such as reducing the similarity between the text and pretraining data.
7 Conclusion
In this paper, we evaluate the extent of GPT-4’s verbatim memory for fictional works and its impact on downstream tasks. We also propose methods to mitigate this impact and encourage the model to use reasoning. We draw several key conclusions: (1) The model always tends to first rely on verbatim memory when performing character understanding tasks, and it consistently benefits from this approach. (2) When the predicted text differs significantly from the pretraining data, the model is more likely to use reasoning to attempt to compensate.
8 Limitations
The specific training data of GPT-4o is not publicly disclosed by OpenAI, making it fundamentally unknowable. Our work employs probabilistic inference to estimate the model’s familiarity with a set of fictional works. However, whether these works are explicitly included in the training data remains unanswerable. Nevertheless, even mere familiarity can suffice to contaminate test data, posing challenges to fair evaluation.
Our study can be extended by future research in a number of ways. First, though we ground and motivate our work through “verbatim” and “gist” memory, we note that any apparent “reasoning” of AI systems may differ fundamentally from those of humans, and we did not pursue human studies to examine this in detail. Further investigation is needed to explore this topic.
Second, the relationship between memorization and reasoning remains a subject of debate in cognitive science. Future studies may uncover insights into this relationship from entirely new perspectives.
Finally, our study was centered on select, recent (but not contemporaneous) majority Western popular culture references. Given their popularity, this selection likely increases the chances that LLMs have been pretrained on more instances or language discussing those references (vs. contemporaneous references, or references from non-Western cultures, where overall occurrences in training data may be lower); it is possible that this exaggerated some of the effects we observed.
9 Ethics
Our study probes the degree to which GPT-4o have memorized fictional works, and its impact on character understanding tasks. This work uses the OpenAI API on experiments above, and at no point do we access, or attempt to access, the true training data behind these models, or any underlying components of the systems. Risks The several datasets and fictional works scripts in our experiment are sourced from publicly available sources. However, we cannot guarantee that they are devoid of socially harmful or toxic language. We use ChatGPT  [^8] https://chatgpt.com/ to correct grammatical errors in this paper.
Appendix A Example of Script-based Character Understanding Task
In Table 8 we show a full example of the character understanding task. In this example, we have automatically replaced the individual character names with placeholders like P0 and P1.
Input P0: Hey, sorry about that P1: No, we’re sorry. We never should have been comparing relationships in the first place. P2: Why? We won. You know, I say, next, we take on Koothrappali and his dog. Really give ourselves a challenge. P3: I just want to say one more thing about this. Just because Penny and I are very different people does not mean that we’re a bad couple. P2: The answer is one simple test away. Hmm? You know, it’s like when I thought there was a possum in my closet. Did I sit around wondering? No, I sent Leonard in with a pointy stick and a bag. P3: I killed his Chewbacca slippers. P0: Let’s just take the test. P3: No, no, no, I don’t want to. P0: Oh, well, ’cause you know we’re gonna do bad. P3: Because it doesn’t matter. I don’t care if we’re a ten or a two. P2: Or a one. A one is possible. P3: Marriage is scary. You’re scared, I’m scared. But it doesn’t make me not want to do it. It, it just makes me want to hold your hand and do it with you. P0: Leonard. P1: It makes me so happy if you said things like that. P2: We got an eight-point-two. Trust me, you’re happy.
Label P0: Penny, P1: Amy, P2: Sheldon, P3: Leonard
Table 8: An example from character guessing task.
Appendix B Example for Verbatim Memory
B.1 Source Prediction
Input Tell me the source for following script:Penny: Hey, sorry about that Amy: No, we’re sorry. We never should have been comparing relationships in the first place. Sheldon: Why? We won. You know, I say, next, we take on Koothrappali and his dog. Really give ourselves a challenge. Leonard: I just want to say one more thing about this. Just because Penny and I are very different people does not mean that we’re a bad couple.
Source The Big Bang Theory
Table 9: An example for source prediction task.
Appendix C Self-generated Character Description
For the self-generated character description, we use following prompts to ask ChatGPT to generate a brief character summarization for each character. The following block-quoted text shows the output we got; importantly, it demonstrates the level of deep knowledge about each of these characters that is baked into the model itself.
Input: Give me a detailed character description for all main characters in the TV show Friends.
Output: Rachel Green (Jennifer Aniston):
Background: Initially introduced as a runaway bride, Rachel starts as a waitress at Central Perk and evolves into a successful fashion executive.
Character Traits: Fashion-conscious, independent, initially somewhat spoiled, but grows into a more self-sufficient and career-focused woman.
Romantic Arc: Has an iconic on-again, off-again relationship with Ross throughout the series.
Ross Geller (David Schwimmer):
Profession: Paleontologist working at a museum.
Character Traits: Intelligent, nerdy, and somewhat socially awkward. Ross has a complicated love life and is often portrayed as the "smart" but emotionally vulnerable member of the group.
Romantic Arc: His tumultuous relationships, particularly with Rachel, are central to the series.
Monica Geller (Courteney Cox):
Profession: Chef and later a restaurant owner.
Character Traits: Competitive, organized, and slightly obsessive about cleanliness. Monica is Ross’s sister and becomes the motherly figure of the group.
Romantic Arc: Her relationship with Chandler Bing becomes a major storyline.
Chandler Bing (Matthew Perry):
Profession: Holds an office job with a mysterious title related to statistical analysis.
Character Traits: Known for his sarcastic humor, commitment issues, and use of humor as a defense mechanism. Evolves into a more mature and emotionally secure character.
Romantic Arc: Develops a significant romantic relationship with Monica.
Joey Tribbiani (Matt LeBlanc):
Profession: Aspiring actor throughout the series.
Character Traits: Easygoing, simple-minded, and often portrayed as the least academically inclined. Known for his catchphrase "How you doin’?" and his love for food.
Romantic Arc: Has numerous short-term relationships, and his genuine friendship with Chandler is a key dynamic.
Phoebe Buffay (Lisa Kudrow):
Profession: Eccentric musician with a unique and mysterious past.
Character Traits: Quirky, free-spirited, and spiritually inclined. Phoebe often provides offbeat humor and memorable songs.
Romantic Arc: Her romantic life includes unique relationships, and her eccentricity adds a distinctive element to the group.
These six characters form the core group of friends in the series, and their individual personalities, relationships, and growth over the ten seasons contribute to the show’s humor, heart, and enduring popularity.
Appendix D Prompts for GPT4
for personality traits extraction ’you are a psychology research assistant designed to help analysis character personalities according to conversations.’ conversation ’Here ends the conversation.Give me a character description for each main character according to this conversation.’
Appendix E Name Replacement and Character Comparison
E.1 Cross-Language Uncommon Name Replacement
With gender swap:[Bojing, Cuixia, Jingjing, Yunsheng, Meilin, Yusong] Without gender swap:[Cuixia, Bojing,Yunsheng,Jingjing,Yusong, Meilin]
Task name Before replacement After replacement
TVShow Monica, Joey, Chandler, Phoebe, Ross, Rachel Bojing, Cuixia, Jingjing, Yunsheng, Meilin, Yusong
CSI Grissom, Catherine, Nick, Sara, Brass Jingbo, Jingwen, Huaqiang, Lihui, Weimin
PERSONET Villefort, The Count, Dantès, Caderousse, Danglars, Albert, Mercédès, Elders, Abbé Busoni, Eugénie, Faria priest, Fernand, Franz, Faria, Morre, Monte Cristo Weina, Nüxia, Danqing, Cailing, Danyu, Ailan, Mengzhe, Lianhua, Busu, Yujun, Fanli, Feiyan, Fangqing, Fangli, Moya, Mingxia
Table 10: Besides the main characters, there are many names of minor characters and passersby appearing in the tasks. In each task, these secondary and passerby names can be dozens of times more than the number of main characters. Replacing these names is complex but not impossible (e.g., with the help of LLMs). However, to align with the TV show approach of only replacing the names of main characters, we only replaced the main characters in all tasks. From the results, it is clear that even replacing only the main characters has had a significant effect.
replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
in the conversation with
[Bojing, Cuixia, Jingjing, Yunsheng, Meilin, Yusong].
The comparison before and after name substitution
The scene overall reflects a group dynamic characteristic of close friends who are comfortable jesting with each other, while also being there in times of crisis or emotional turmoil.
1.Monica: Responsible, caring, and organized. Assertive and confident in her actions.
1. Bojing: Bojing comes across as practical, level-headed and caring. He often acts as the voice of reason for his friends, attempting to mediate, clarify, and console in various situations. His attempts to play down his date suggest he is a private person who doesn’t enjoy sharing intimate details of his life.
2. Joey: Energetic, extroverted, and casual. Lacks the sensitivity of others’ feelings at times but genuinely care about friends.
2. Cuixia: Cuixia seems like a lively, fun-loving character. However, she sometimes shows a more cynical side, quick to suspect something might be wrong with Bojing’s date and suggesting a strip joint as a solution to Meilin’s woes.
3. Chandler: Witty and self-deprecating with an approachable sense of humor. Exhibits insecurity and anxiety in his dialogues, making references to uncomfortable situations and questioning his own actions.
3. Jingjing: Jingjing can be open and candid about his thoughts, even if they seem inappropriate or unusual. He’s also humorously self-aware, undercutting his moments of honesty with reminders that he might be oversharing.
4. Phoebe: Quirky, eccentric, and a free spirit. Her train of thought tends to lean toward the unusual and bizarre. However, she is also compassionate and caring.
4. Yunsheng: Appears offbeat and unusual, suggesting the eating chalk anecdote about his past relationship. He also believes in new-age concepts like auras, showing a more spiritual side.
5. Ross: Insecure and somewhat neurotic and vulnerable. His behavior is indicative of someone who is going through emotional turmoil.
5. Meilin: Exhibits vulnerability and emotional turmoil, especially regarding his recent divorce. He seems to be fluctuating between hurt, anger, and longing for his past relationship.
6. Rachel: Spontaneous and open to change, she takes risks and is adaptable. She is also reliant on her relationships with others, signifying her dependency and need for support.
6. Yusong: Yusong presents as impulsive, high-strung and somewhat comical in moments of panic. Fleeing her wedding because of a sudden realization shows she can make drastic decisions based on her emotions.
E.2 Same-Language Replacement
With Gender-Matched replacement, we replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
in the conversation with
[Sally, Andrew, Ethan, Olivia, Benjamin, Lauren].
With Gender Swap as well, we got [Andrew, Sally,Olivia, Ethan, Lauren,Benjamin].
E.3 Anonymized Speaker Replacement
replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
in the conversation with
[Speaker 0, Speaker 1, Speaker 2, Speaker 3, Speaker 4, Speaker 5]. LLM summarized character description
We use the prompt of "you are a psychology research assistant designed to help analysis character personalities according to conversations……..Here ends the conversation.Give me a character description for each main character according to this conversation." " and we get such result from TVshow Friends scene1.
E.4 Control Test
Example Here is a dialogue between 2 people: Sheldon : (now lifts her head) phoebe!! Amy : you know it’s me? Sheldon : for like a half an hour! man, you can lie about sweden! Amy : how can you come here? Sheldon : how could you not tell me you worked here? Amy : i don’t have to tell you everything!P2: We got an eight-point-two. Trust me, you’re happy.
Prompt The following is a dialogue between the same two people. Please identify who they are:
P0 : (singing) ïpan stripan, glupi glabi! änd that is the swedish national anthem! thank you for asking! (looks annoyed) P1 : wow, ikea… what a rich culture. uhm, you know what? i have a friend who is a masseuse. P0 : oh, ja! ja! P1 : yah! she’s… uhm… not very good though… (phoebe looks devastated)
Table 11: An example for the control test, where in the example,the script is from Friends but speakers are annoted as Amy and Sheldon(from The Big Bang Theory).
Appendix F Verbatim avoid Reasoning
Figure 3: In a personality understanding task where the scenario involves ’Joey returned the lost wallet to its rightful owner, Monica,’ the question posed might be, ’According to this action, his personality should be: ?’ The human reasoning process would draw on contextual evidence, concluding that returning lost property demonstrates honesty. In contrast, the LLM might conclude, ’Joey and Monica are from the TV show Friends, and Joey is characterized as an honest character in this show,’ thus relying heavily on memorized information rather than genuine reasoning.
Appendix G Benchmark Datasets
The six selected benchmarks cover a range of character-centric tasks, designed to test different aspects of character understanding(Table 1):
  • TVShow Guessing: Based on the TVSHOWGUESS dataset, which contains scripts from five popular TV series: Friends, The Big Bang Theory, The Office, Frasier, and Gilmore Girls Sang et al. (2022c), this is a multi-choice classification task. The dialogue is anonymized using placeholder identifiers (e.g., P1, P2), and the model must predict the correct speaker for each line, testing its ability to identify characters based on context.
  • Screenplay: Using screenplays from nine popular movies, such as Avengers: Endgame (2019), Dead Poets Society (1989), John Wick (2014), and The Shawshank Redemption (1994) Baruah and Narayanan (2023), this is a coreference resolution task. The model is tasked with linking character mentions throughout the text, forming coreference clusters and resolving ambiguities in character references.
  • Personality Understanding: Using the PERSONET dataset, which comprises approximately 32,000 samples from 33 classic literature books Yu et al. (2023b), this task involves predicting the most likely personality trait depicted in a given text snippet, along with its preceding context. Formulated as a multi-choice problem with distractor traits, this task evaluates the model’s ability to infer character personalities from descriptive text.
  • CSI Role Detection: Based on 39 episodes of the TV series CSI Frermann et al. (2018), this sequence labeling task requires the model to identify mentions of the perpetrator in dialogues and scene descriptions, focusing on role detection within crime-related narratives.
  • FriendsQA: Derived from 1,222 scenes from the first four seasons of Friends Yang and Choi (2019), this dataset supports an open-domain question answering task. The model is given a dialogue excerpt and a manually crafted question and must extract the specific span in the dialogue that answers the question, testing its comprehension of dialogue context.
  • SUMMSCREEN: Built from transcript-recap pairs of 91 TV series collected from community sources such as The TV MegaSite, Inc. (TMS) and ForeverDreaming (FD) Chen et al. (2022), this dataset supports a summarization task. The model is required to generate a coherent summary from dialogue utterances and scene descriptions, capturing key plot points while excluding irrelevant details.