### v4

# Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation

Abstract
Recently, Large Language Models (LLMs) have shown impressive performance in various character understanding tasks. However, the extensive pre-training corpora used by LLMs raise concerns that they may rely on memorizing popular fictional works rather than genuinely understanding and reasoning about them. Prior work has addressed memorization in LLM reasoning tasks, but these approaches do not directly apply to character understanding. To address this, we leverage the cognitive science concepts of "gist memory" and "verbatim memory", where verbatim memory corresponds to exact recall, while gist memory captures underlying meaning. We design a hard-soft intervention framework, where the hard setting disrupts memorization by perturbing key entities, while the soft setting maintains the original text but guides the model toward reasoning. Our experiments across six character understanding tasks with GPT-4o, LLaMA3.3-70B, and DeepSeek V3 demonstrate that reducing verbatim memory reliance shifts model behavior toward reasoning. Our method leads to a performance drop of up to 45.8% when models can no longer depend on memorization, highlighting the extent to which existing benchmarks inflate performance through data contamination. Additionally, we show that prompting techniques can systematically alter a model’s response strategy, making it more dependent on either memorization or reasoning.
Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation
Yuxuan Jiang , Francis Ferraro Department of Computer Science and Electrical Engineering University of Maryland, Baltimore County yuxuanj1@umbc.edu
1 Introduction
Character understanding requires the models to grasp the characters’ personalities and traits while reading a story, which aims to evaluate the models’ capability to comprehend personalities, goals, roles, and relationships of various fictional characters Yu et al. (2023a); Vallurupalli et al. (2024); Shen et al. (2024); Brahman et al. (2021). However, the impressive performance of LLMs on such benchmarks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023) raises the question: does it reflect genuine understanding or merely strong memorization Chang et al. (2023a)? For example, in the TV show speaker guessing task shown in Figure  1, the model is expected to identify Sheldon based on its memory of the character from previous episodes (e.g., preparing lessons for Howard). However, if the model simply memorizes the exact words, it could make predictions based on a very short sentence by directly extracting the speaker information from it.
To truly stimulate a model’s character understanding ability, it is crucial to understand how the model may be generating its responses. This can be especially challenging given the extensive and opaque pre-training corpora used by LLMs, making it particularly severe for character understanding tasks, as many benchmarks are derived from well-known and widely consumed fiction or scripts Sang et al. (2022a). We hypothesize that this can lead to a form of memorization called “verbatim memory,” which encodes precise details and remembers every word exactly Huang et al. (2024); Ippolito et al. (2022)—rather than reasoning over recalled overarching patterns (what is referred to as “gist memory” Reyna and Brainerd (1998)).
Figure 1: This figure illustrates the distinct mechanisms of gist and verbatim memory processes in the speaker-guessing task for a script. Verbatim memory (bottom right) aims to recall as much specific information as possible, including who said each sentence. If the memorization is not precise enough, the model fails to provide the correct answer. In contrast, gist memory (top right) focuses on extracting core information, such as character relationships, personalities, and key events. By relying on gist memory, the model is encouraged to engage in reasoning rather than purely relying on memorization.
Motivated by the concept of gist memory, we propose a prompt-based approach to mitigate model reliance on verbatim memorization in character understanding tasks. We evaluate our method on GPT-4o, LLaMA3-70B, and DeepSeek V3, introducing a hard-soft setting framework: the hard setting blocks memorization by replacing character names, while the soft setting retains the original text but employs gist-based prompts to encourage reasoning over recall. Our ablation study confirms the effectiveness of both interventions.
Building on these findings, we extend our approach to five character understanding tasks, designing task-specific gist prompts and evaluating models under both settings. Our results indicate that the Gist+NR approach best reduces memorization dependency.
In summary, we outline our study and key contributions as follows:
  1. 1.
    Character understanding tasks heavily rely on verbatim memorization. Across six benchmark tasks, all evaluated LLMs show significant performance drops when explicit memorization is disrupted, confirming that these tasks often assess recall rather than true reasoning (Table 6.2).
  2. 2.
    Gist-based prompting effectively reduces memorization dependency. Introducing gist-based prompts systematically lowers reliance on verbatim recall across all tasks, encouraging models to use reasoning instead of mechanical memorization (Section 6).
  3. 3.
    Verbatim and gist-based prompting modulate model response strategies. Models adjust their response strategies based on the prompting style—verbatim-based prompts reinforce reliance on memorization, while gist-based prompts encourage reasoning-driven answers (Section 4.3).
2 Related Work
2.1 Character Understanding and its evolution with LLM
Character understanding has made significant contributions to advancing story comprehension in psychological, literary, and educational research  Bower and Morrow (1990); McKee (2010); Paris and Paris (2003). This multifaceted field encompasses a variety of tasks, including character recognition, coreference resolution, summarization, role detection, goal understanding, personality analysis, and question answering Sang et al. (2022b); Bamman et al. (2020); Chen et al. (2022); Frermann et al. (2018); Vallurupalli et al. (2024); Yu et al. (2023b). Recent studies have shown promising performance by LLMs on character understanding tasks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023). However, these studies often overlook the potential influence of memorization on model performance. Fictional works, including movie and TV show scripts, famous novels, and news texts, serve as primary sources for narrative understanding Sang et al. (2022a); Tapaswi et al. (2016); Caselli and Vossen (2017). These texts, much like copyrighted books Chang et al. (2023a), are prone to strong memorization. Naturally, this raises curiosity about how memorization affects downstream tasks, particularly character understanding. While previous research has demonstrated the memorization capabilities of LLMs, our work further explores its specific impact on character understanding and proposes an approach to mitigate this effect.
2.2 LLMs’ Memorization and Reasoning
Among LLMs’ various capabilities Li et al. (2024c, e); Zhang et al. (2024); Li et al. (2024d), the reasoning abilities Tong et al. (2023); Yang et al. (2023); Xiong et al. (2024b); Tong et al. (2024); Xiong et al. (2024a) of LLMs Li et al. (2024b); Tan et al. (2024) have garnered significant attention, but their memorization has also been shown to play a crucial role in influencing benchmark reasoning performance Zhou et al. (2023); Xie et al. (2024a); Chen et al. (2024); Li et al. (2024a); Biderman et al. (2024). The influence of memorization is complex: data contamination can inflate performance metrics Khandelwal et al. (2020); Blevins and Zettlemoyer (2022); Carlini et al. (2023), while knowledge conflict can reduce them Su et al. (2024); Wang et al. (2024); Xie et al. (2024b); Xu et al. (2024).
While prior work has focused on detecting and mitigating memorization effects in reasoning tasks such as numerical, commonsense, and logical reasoning Xie et al. (2024a); Sainz et al. (2023); Li et al. (2024f), character understanding tasks introduce additional challenges due to their diverse formats and varying memory demands. Unlike structured reasoning tasks where excessive memorization is clearly undesirable, character understanding often requires a nuanced balance between recall and inference.
2.3 Verbatim and Gist Memory in LLM Reasoning
Cognitive science distinguishes between verbatim memory, which retains exact details, and gist memory, which captures the essential meaning Brainerd and Reyna (2002); Reyna and Brainerd (1998). For example, verbatim memory would store a sentence like "The black-and-white cat sat on a soft, red mat in the living room" with all details intact, while gist memory would reduce it to "A cat was sitting indoors," focusing on the core idea while discarding specifics. Human reasoning primarily relies on gist memory, as the two types of memory function independently rather than hierarchically Brainerd and Reyna (1992).
In LLM research, memorization is often equated with verbatim memory, where models retrieve and reproduce exact strings from pretraining data Carlini et al. (2023). A common mitigation strategy involves curating test samples that the model has never seen before Xie et al. (2024a). However, in character understanding tasks, acquiring obscure or newly released datasets can be costly and difficult to control in terms of quality. Instead, our work proposes leveraging gist memory to reduce reliance on precise memorization, encouraging LLMs to generalize character understanding beyond string-matching. By shifting the focus toward gist memory, we aim to align LLM character analysis with human-like reasoning rather than dataset recall.
3 Data and Experiment Setup
In our work, we utilize two distinct datasets to evaluate LLMs on: (1) their reasoning capabilities in character understanding under gist memory and (2) their memorization of original source texts.
3.1 Benchmark Datasets for Character Understanding Tasks
To evaluate the interplay between memorization and reasoning in character understanding tasks, we employ six publicly available datasets covering diverse task types. These datasets span different narrative mediums, including TV shows, movies, and novels, ensuring diversity in both content and character representation. The task formulations and dataset compositions are detailed below, with an overview provided in Table 1. Further task details are available in  6.1.
Datasets
We evaluate character understanding using six publicly available datasets spanning diverse sources, including TV shows, movies, and novels. For coreference resolution, we use the latest release of the MovieCoref dataset  [^1] https://github.com/usc-sail/mica-character-coref, which provides annotated coreference chains in movie scripts. For character guess, we adopt TVSHOWGUESS  [^2] https://github.com/YisiSang/TVSHOWGUESS, where models must identify the speaker of a given dialogue. For personality understanding, we use PERSONET  [^3] https://github.com/Gorov/personet_acl23, a dataset derived from literary sources with personality trait annotations. For role detection, we employ CSI  [^4] https://github.com/EdinburghNLP/csi-corpus, which focuses on identifying key characters in crime narratives. For open-domain question answering, we evaluate models on FriendsQA  [^5] https://github.com/emorynlp/FriendsQA, where questions are answered by extracting spans from TV show dialogues. Lastly, for summarization, we use SummScreen  [^6] https://github.com/mingdachen/SummScreen, which pairs TV show transcripts with corresponding summaries. We follow the original dataset authors’ preprocessing, development, and test splits to ensure comparability. Further dataset details are provided in Appendix G.
Evaluation Metrics
We followed the original evaluation metrics used in the respective tasks, shown in Table 1.
Task Name Task Type Data Source Metrics
TVShow Guess Character Guessing TV Shows Exact Match
FriendsQA Question Answering TV Shows Exact Match
Screenplay Character Coreference Movies F1 Score
SummScreen Text Summarization TV Shows ROUGE
CSI Role Detection TV Shows F1 Score
PERSONET Personality Understanding Novels Exact Match
Table 1: Overview of benchmark tasks, including their type, data sources, and evaluation metrics.
3.2 Fictional Works Memorization Benchmark Dataset
This dataset is designed to evaluate LLMs’ memorization of fictional scripts by assessing their ability to recall verbatim details from diverse narrative sources.
Benchmark Sources
The six benchmark datasets described in Section 6.1 are derived from a diverse set of sources, including 91 TV shows, 12 movies, and 33 classic literary works.
Dataset Construction
To ensure broad coverage across different narrative mediums, we sample 200 script segments from 40 selected sources, comprising 11 TV series, 9 movies, and 20 novels. Each source contributes 5 excerpts, with an average segment length of 312 words.
Segment Format
Each excerpt follows the format: TV Show Title; Scene, where the scene consists of a mixture of dialogue and narration extracted from the original script.
3.3 Models
we evaluate both open-source and closed-source models, leveraging their APIs for inference. Our evaluation includes three frontier LLMs: the open-source Llama 3.3-70B AI (2024) and DeepSeek V3 Liu et al. (2024), as well as the closed-source GPT-4o OpenAI (2023).
4 Soft Setting: Memorization-Controlled Prompting in Character Guessing
First, we investigate whether prompting strategies can subtly influence LLMs’ response patterns in a soft setting, where the model is not explicitly restricted from using memorization but is instead encouraged to rely more on reasoning over direct retrieval.
4.1 Character Guessing Task Format
Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where each utterance $U_{i}$ is paired with a speaker label $S_{i}$. In the Character Guessing Task, the model is given a dialogue excerpt without speaker annotations and must predict the correct speaker for each utterance, $f_{\text{guess}}(U_{i})\rightarrow S_{i}$. We follow the original format and setting Sang et al. (2022b), using the TVShow Guess Dataset described in Section 3.1. This task is designed to prompt the model to make guesses based on character relationships, key events, and personality traits within the text. However, its format closely resembles the name cloze tests Chang et al. (2023b) recently used to evaluate LLM memorization.
4.2 Prompting Strategies
We design two distinct prompting conditions to examine how LLMs rely on different memory mechanisms when identifying characters. Verbatim prompting encourages mechanical memorization, where the model recalls exact substrings from its training data and reconstructs missing parts without reasoning. In contrast, gist prompting focuses on extracting key information and performing memory-based matching. For character understanding tasks, the most relevant gist cues typically include character relationships, significant events, and personality traits.
All prompts share the same base instruction: "The following is a dialogue. Please identify who they are. …"
The two variations are as follows:
Verbatim Prompt: "…Do this by direct memory retrieval—do not infer based on traits, interactions, or context. Recall names exactly from past dialogue knowledge without explanations or reasoning."
Gist Prompt: "…Do this by analyzing relationships, key events, and personality traits to match them with known characters. Do not directly retrieve from memory who said the sentences."
4.3 Result and Analysis
Figure 2 visualizes the performance of the three models on the Character Guessing task under different prompting conditions. The performance trends remain consistent across models, and our key observations are as follows:
  • Baseline Performance: The baseline condition reflects the overall performance of the models, with GPT-4o demonstrating a clear lead, while Llama 3.3-70B and DeepSeek V3 exhibit similar performance.
  • Memorization Reliance: All three models show that the Verbatim condition performs nearly identically to the baseline, supporting our hypothesis that models primarily rely on memorization when answering the Character Guessing task.
  • Effect of Gist Prompting: The Gist condition results in a non-trivial performance drop, suggesting that our prompt effectively reduces the models’ reliance on memorization when generating responses.
Figure 2: The bar chart compares the performance of three LLMs under three prompting settings. Across all three models, the performance of the Verbatim and Baseline conditions is highly similar, confirming our hypothesis that the models rely on memorization when answering questions. The Gist condition, however, shows a non-trivial decline in performance, indicating that the models struggle when they cannot fully rely on memorization to generate responses.
5 Hard Setting: Memorization-Controlled Prompting under Name Replacement
Furthermore, we aim to validate whether our prompting strategies can genuinely alter the model’s reasoning patterns under a hard setting, where memorization is explicitly restricted. Previous research has suggested that name replacement may impact model performance Yu et al. (2022). In this section, we confirm this hypothesis by demonstrating that name replacement significantly affects the model’s memorization ability and quantifying its impact. Then, we examine how prompting strategies influence LLMs’ reliance on memorization versus reasoning when memorization is constrained by name replacement.
5.1 Name Replacement Strategies and Their Impact on Memorization
Name Replacement Strategies
Previous studies Yu et al. (2022) have primarily employed Same-Cultural Name Replacement, a naive and intuitive approach that substitutes character names with culturally consistent alternatives. To systematically analyze the impact of name replacement on model performance, we introduce three distinct levels of name modifications:
  • Name Masking: Replacing all character names with placeholders (e.g., Sheldon → P0, Amy → P1).
  • Cross-Cultural Name Replacement: Substituting names with those from a different cultural background while maintaining gender consistency (e.g., SheldonJie Zhang, AmyXiaoling Wang). This tests whether models rely on specific name distributions for character inference.
  • Same-Cultural Name Replacement: Replacing names with other culturally consistent alternatives (e.g., SheldonAndrew, AmySally). This retains cultural familiarity while disrupting direct memorization.
Memorization Evaluation Task Format
We design the Source Identification Task where the model is given a dialogue excerpt without source annotations and must predict the correct origin of the text. Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where each utterance $U_{i}$ is part of a larger script from a specific source. The task requires the model to determine the correct source, $f_{\text{source}}(D)\rightarrow S$, where $S$ denotes the source work. We use the Fictional Works Memorization Benchmark Dataset described in Section 3.2.
Impact of Name Replacement on Memorization.
According to Figure 3, our results demonstrate that different name replacement strategies have a significant impact on the memorization ability of LLMs. Among them, Cross-Cultural Name Replacement consistently induces the strongest forgetting effect across all models, indicating that altering names to those from a different cultural background disrupts memorized associations more effectively than other replacement methods.
Figure 3: The heatmap presents accuracy percentages for source identification across three LLMs under different name replacement conditions. All three models demonstrate strong memorization of the script. The Cross-Culture Name Replacement method exhibits a significantly strong and the most pronounced memorization-weakening effect across all three models.
5.2 Memorization-Controlled Prompting under Name Replacement
Different name replacement strategies have shown a significant impact on the memorization ability of LLMs. Building on this, we apply the name replacement strategies introduced in Section 5.1 to evaluate the Character Guessing task using the different prompting strategies discussed in Section 4.2.
Under the Name Replacement condition, we provide brief character descriptions generated by GPT-4, aiming to enable the model to complete the Character Guessing task through character matching. The descriptions used are provided in the Appendix D. To assess the actual effectiveness of these descriptions, we conduct an ablation study.
Figure 4: The line chart shows the decline in memorization across different settings, illustrating that as memorization weakens, overall model performance also decreases. The bar chart presents the effect of different prompting strategies: under Name Replacement conditions, Verbatim prompting consistently outperforms the baseline by reinforcing memorization, whereas Gist prompting further reduces reliance on memorization, leading to the lowest accuracy.
Results and Key Findings
We present the results in Figure 4 and summarize our main findings as follows:
  1. 1.
    Strong Correlation Between Memorization and Performance. Model performance aligns closely with memorization levels, indicating a strong dependence on memorization for accurate predictions.
  2. 2.
    Minimal Impact of Character Descriptions . As an ablation study, we evaluate the effect of character descriptions under the Origin Script condition. Our results show that incorporating character descriptions does not significantly affect model performance, ruling out potential interference from inaccurate or misleading descriptions.
  3. 3.
    Verbatim Prompting Enhances Memorization. Across all three name replacement strategies, Verbatim prompting consistently improves model performance over the baseline, demonstrating that prompting can reinforce memorization and consequently boost accuracy.
  4. 4.
    Gist Prompting Reduces Reliance on Memorization. Performance under the Gist condition declines compared to the baseline, confirming that our Gist prompt effectively reduces the model’s reliance on memorization.
  5. 5.
    Name Replacement + Gist Prompting as an Effective Evaluation Strategy. We conclude that combining Cross-Culture Name Replacement with Gist prompting provides a reasonable and effective method for evaluating a model’s true capabilities while minimizing reliance on memorization.
6 Evaluating Character Understanding Tasks Under Memorization Constraints
In this section, we evaluate different types of character understanding tasks using the Gist Prompting with Cross Culture Name Replacement approach. We first introduce the design principles for Gist Prompts across various task types, followed by an analysis of the results obtained from three different LLMs.
6.1 Benchmark Tasks Description
We define the five character understanding tasks being benchmarked and provide an illustrative example.
Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where $U_{i}$ is an utterance in the conversation. Each utterance $U_{i}$ is associated with a speaker label $S_{i}$, and additional annotations depending on the task.
  • Coreference Resolution: Cluster character mentions $C=\{c_{1},c_{2},...,c_{n}\}$ into groups referring to the same entity, $f_{\text{coref}}(C)\rightarrow\{\text{clusters of }c_{i}\}$.
  • Personality Understanding: Classify a character’s personality trait $P$ given dialogue context $D$, $f_{\text{personality}}(D,S_{i})\rightarrow P$.
  • Role Detection: Determine whether a character plays a key role (e.g., perpetrator in a crime narrative), $f_{\text{role}}(S_{i})\rightarrow\{0,1\}$.
  • Open-Domain Question Answering: Extract the span $A$ in $D$ that answers a given question $Q$, $f_{\text{QA}}(D,Q)\rightarrow A$.
  • Summarization: Generate a summary $S$ that captures key events and relationships, $f_{\text{summ}}(D)\rightarrow S$.
Example Dialogue
To illustrate these tasks, consider the following conversation:
P0: "You’re still going to have to convince a jury that I killed two strangers for no reason."
Grissom doesn’t look worried. He takes his gloves off and puts them on the table.
P1: "You ever been to the theater, Peter? There’s a play called Six Degrees of Separation."
The correct speaker labels are P0 → Peter Berglund, P1 → Grissom. For all tasks except Character Guess, the placeholders P0 and P1 are replaced with their original names, Peter Berglund and Grissom. Table 2 illustrates how this dialogue segment is utilized across different tasks.
Task Example Question and Expected Output
Coreference Resolution "Which entity does ’his’ refer to?" → "Grissom"
Personality Understanding "What trait best describes P1?" → Analytical
Role Detection "Identify all entities referring to the criminal." → Peter Berglund
Question Answering "What play does Grissom mention?" → "Six Degrees of Separation"
Summarization "Generate a summary." → "Peter discusses convincing a jury, while Grissom references a play."
Table 2: Examples of different character understanding tasks applied to the same dialogue.
6.2 Gist Prompts Across Character Understanding Tasks
General Guidelines for Gist Prompting
  • Encourage Structural Reasoning: Guide models to infer character identity and relationships from discourse patterns rather than relying on explicit name associations.
  • Eliminate Memorization Bias: Ensure prompts emphasize evidence-based inference rather than knowledge retrieval.
  • Preserve Narrative Coherence: In summarization and question answering, focus on maintaining logical flow rather than phrase-matching.
Detailed gist prompts are included in Appendix H.
{NiceTabular}
l|ccc|ccc|ccc Tasks GPT-4o LLaMA3.3-70B DeepSeek V3
Origin NR NR+GIST Origin NR NR+GIST Origin NR NR+GIST
Tvshow Guess 78.2 32.0 $\downarrow$46.2 31.5 $\downarrow$0.5 61.4 31.9 $\downarrow$29.5 27.8 $\downarrow$4.1 60.8 29.7 $\downarrow$31.1 26.7 $\downarrow$3.0
Coreference 58.3 49.6 $\downarrow$8.7 47.5 $\downarrow$2.1 56.7 48.3 $\downarrow$8.4 48.1 $\downarrow$0.2 47.4 34.6 $\downarrow$12.8 32.3 $\downarrow$2.3
FriendsQA 44.3 38.6 $\downarrow$5.7 38.6 $\uparrow$0.0 45.7 42.9 $\downarrow$2.8 44.3 $\uparrow$1.4 42.8 39.4 $\downarrow$3.4 34.3 $\downarrow$5.1
ScreenSum 37.5 28.7 $\downarrow$8.8 23.1 $\downarrow$5.6 40.1 25.9 $\downarrow$14.2 21.3 $\downarrow$4.6 35.6 19.3 $\downarrow$16.3 17.1 $\downarrow$2.2
CSI Role Extract 50.1 32.9 $\downarrow$17.2 31.7 $\downarrow$1.2 48.3 34.7 $\downarrow$13.6 32.7 $\downarrow$2.0 45.1 29.8 $\downarrow$15.3 23.7 $\downarrow$6.1
PERSONET 56.0 52.0 $\downarrow$4.0 54.0 $\uparrow$2.0 62.0 46.0 $\downarrow$16.0 44.0 $\downarrow$2.0 44.0 40.0 $\downarrow$4.0 38.0 $\downarrow$2.0
Table 3: Ablation study examining the impact of name replacement and the GIST method on model performance across six NLP tasks.The subscripted values in the NR column indicate the performance drop relative to the original script, while the subscripted values in the NR+GIST column show the performance change compared to the Name Replace condition.
6.3 Results and Key Findings
Our main results and findings are summarized as follows:
  1. 1.
    Significant Performance Drop with Name Replacement: Across all tasks, we observe a substantial decline in performance when applying the name replacement (NR) method in Table 6.2. This result indicates that reducing explicit memorization has a strong impact on model performance, confirming our hypothesis that existing character understanding tasks rely heavily on memorization. Moreover, it highlights the potential inflation of benchmark scores under standard evaluation settings.
  2. 2.
    Further Degradation with Gist Prompting: The additional performance drop induced by Gist prompting demonstrates its effectiveness across various character understanding tasks. This suggests that Gist prompting can further encourage models to rely less on memorization, reinforcing our findings on the role of memorization in these tasks.
  3. 3.
    Memorization Can Be Detrimental to Character Understanding: Interestingly, we observe performance gains in certain cases under the combined NR+Gist setting. Our error case analysis reveals that these improvements stem from resolving knowledge conflicts—instances where the original memorization-based responses misled the model, whereas reducing memorization mitigated such errors, leading to better predictions.
  4. 4.
    Consistent Trends Across Models: As shown in Figure 5, despite variations in absolute performance, different models exhibit similar trends in performance degradation across all tasks. This consistency suggests that various task designs interact similarly with memorization, reinforcing the broader implications of our findings.
Figure 5: Performance degradation across LLMs due to the effects of GIST+NR perturbations.
7 Conclusion
In this work, we introduce the concept of gist memory and design prompts for six types of character understanding tasks, enabling models to reduce reliance on memorization and enhance reasoning-based character interpretation. Our approach provides a systematic framework to evaluate the extent to which different character understanding tasks depend on memorization, offering a more accurate assessment of model performance by disentangling genuine reasoning from memorization-driven responses. These findings highlight the limitations of current benchmarks and provide insights for designing more robust character understanding evaluations.
8 Limitations
The specific training data of GPT-4o is not publicly disclosed by OpenAI, making it fundamentally unknowable. Our work employs probabilistic inference to estimate the model’s familiarity with a set of fictional works. However, whether these works are explicitly included in the training data remains unanswerable. Nevertheless, even mere familiarity can suffice to contaminate test data, posing challenges to fair evaluation.
Our study can be extended by future research in a number of ways. First, though we ground and motivate our work through “verbatim” and “gist” memory, we note that any apparent “reasoning” of AI systems may differ fundamentally from those of humans, and we did not pursue human studies to examine this in detail. Further investigation is needed to explore this topic.
Second, the relationship between memorization and reasoning remains a subject of debate in cognitive science. Future studies may uncover insights into this relationship from entirely new perspectives.
Finally, our study was centered on select, recent (but not contemporaneous) majority Western popular culture references. Given their popularity, this selection likely increases the chances that LLMs have been pretrained on more instances or language discussing those references (vs. contemporaneous references, or references from non-Western cultures, where overall occurrences in training data may be lower); it is possible that this exaggerated some of the effects we observed.
9 Ethics
Our study probes the degree to which GPT-4o have memorized fictional works, and its impact on character understanding tasks. This work uses the OpenAI API on experiments above, and at no point do we access, or attempt to access, the true training data behind these models, or any underlying components of the systems. Risks The several datasets and fictional works scripts in our experiment are sourced from publicly available sources. However, we cannot guarantee that they are devoid of socially harmful or toxic language. We use ChatGPT  [^7] https://chatgpt.com/ to correct grammatical errors in this paper.
Appendix A Example of Script-based Character Understanding Task
In Table 4 we show a full example of the character understanding task. In this example, we have automatically replaced the individual character names with placeholders like P0 and P1.
Input P0: Hey, sorry about that P1: No, we’re sorry. We never should have been comparing relationships in the first place. P2: Why? We won. You know, I say, next, we take on Koothrappali and his dog. Really give ourselves a challenge. P3: I just want to say one more thing about this. Just because Penny and I are very different people does not mean that we’re a bad couple. P2: The answer is one simple test away. Hmm? You know, it’s like when I thought there was a possum in my closet. Did I sit around wondering? No, I sent Leonard in with a pointy stick and a bag. P3: I killed his Chewbacca slippers. P0: Let’s just take the test. P3: No, no, no, I don’t want to. P0: Oh, well, ’cause you know we’re gonna do bad. P3: Because it doesn’t matter. I don’t care if we’re a ten or a two. P2: Or a one. A one is possible. P3: Marriage is scary. You’re scared, I’m scared. But it doesn’t make me not want to do it. It, it just makes me want to hold your hand and do it with you. P0: Leonard. P1: It makes me so happy if you said things like that. P2: We got an eight-point-two. Trust me, you’re happy.
Label P0: Penny, P1: Amy, P2: Sheldon, P3: Leonard
Table 4: An example from character guessing task.
Appendix B Example for Verbatim Memory
B.1 Source Prediction
Input Tell me the source for following script:Penny: Hey, sorry about that Amy: No, we’re sorry. We never should have been comparing relationships in the first place. Sheldon: Why? We won. You know, I say, next, we take on Koothrappali and his dog. Really give ourselves a challenge. Leonard: I just want to say one more thing about this. Just because Penny and I are very different people does not mean that we’re a bad couple.
Source The Big Bang Theory
Table 5: An example for source prediction task.
Appendix C Extended Dataset Analysis
For the six representative character understanding tasks studied in our work, we selected one dataset for each category. However, we also recognize that there are many different benchmark datasets for each type of task. When selecting representative datasets for each category, we focused on choosing those with more diverse sources. Additionally, we supplemented the selection with eight datasets of the same type but derived from different sources in Table 6.
Task Name Task Type Source
Character Mining Character Identification TV Shows
LISCU Character Identification Book Reviews
MovieQA Visual Question Answering Movies
Liter Coreference Character Coreference Novels
HVV Detect Role Detection Movies, News
SYMON Visual Question Answering Movies
CEB Character Representation Novels
Love and Violence Summarization TV Shows
Table 6: Overview of 14 tasks in the character understanding benchmark.
Appendix D Prompts for GPT4
for personality traits extraction ’you are a psychology research assistant designed to help analysis character personalities according to conversations.’ conversation ’Here ends the conversation.Give me a character description for each main character according to this conversation.’ Character Description Here is a character description:Monica: Responsible, caring, and organized. Assertive and confident in her actions.Chandler: Witty and self-deprecating with an approachable sense of humor. Exhibits insecurity and anxiety in his dialogues, making references to uncomfortable situations and questioning his own actions.Rachel: Spontaneous and open to change, she takes risks and is adaptable. She is also reliant on her relationships with others, signifying her dependency and need for support.’ ’Joey: Energetic, extroverted, and casual. Lacks the sensitivity of others feelings at times but genuinely care about friends.Phoebe: Quirky, eccentric, and a free spirit. Her train of thought tends to lean toward the unusual and bizarre. However, she is also compassionate and caring.Ross: Insecure and somewhat neurotic and vulnerable. His behavior is indicative of someone who is going through emotional turmoil.’
Appendix E Name Replacement and Character Comparison
E.1 Cross-Language Uncommon Name Replacement
With gender swap:[Bojing, Cuixia, Jingjing, Yunsheng, Meilin, Yusong] Without gender swap:[Cuixia, Bojing,Yunsheng,Jingjing,Yusong, Meilin]
Task name Before replacement After replacement
TVShow Monica, Joey, Chandler, Phoebe, Ross, Rachel Bojing, Cuixia, Jingjing, Yunsheng, Meilin, Yusong
CSI Grissom, Catherine, Nick, Sara, Brass Jingbo, Jingwen, Huaqiang, Lihui, Weimin
PERSONET Villefort, The Count, Dantès, Caderousse, Danglars, Albert, Mercédès, Elders, Abbé Busoni, Eugénie, Faria priest, Fernand, Franz, Faria, Morre, Monte Cristo Weina, Nüxia, Danqing, Cailing, Danyu, Ailan, Mengzhe, Lianhua, Busu, Yujun, Fanli, Feiyan, Fangqing, Fangli, Moya, Mingxia
Table 7: Besides the main characters, there are many names of minor characters and passersby appearing in the tasks. In each task, these secondary and passerby names can be dozens of times more than the number of main characters. Replacing these names is complex but not impossible (e.g., with the help of LLMs). However, to align with the TV show approach of only replacing the names of main characters, we only replaced the main characters in all tasks. From the results, it is clear that even replacing only the main characters has had a significant effect.
replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
in the conversation with
[Bojing, Cuixia, Jingjing, Yunsheng, Meilin, Yusong].
The comparison before and after name substitution
The scene overall reflects a group dynamic characteristic of close friends who are comfortable jesting with each other, while also being there in times of crisis or emotional turmoil.
1.Monica: Responsible, caring, and organized. Assertive and confident in her actions.
1. Bojing: Bojing comes across as practical, level-headed and caring. He often acts as the voice of reason for his friends, attempting to mediate, clarify, and console in various situations. His attempts to play down his date suggest he is a private person who doesn’t enjoy sharing intimate details of his life.
2. Joey: Energetic, extroverted, and casual. Lacks the sensitivity of others’ feelings at times but genuinely care about friends.
2. Cuixia: Cuixia seems like a lively, fun-loving character. However, she sometimes shows a more cynical side, quick to suspect something might be wrong with Bojing’s date and suggesting a strip joint as a solution to Meilin’s woes.
3. Chandler: Witty and self-deprecating with an approachable sense of humor. Exhibits insecurity and anxiety in his dialogues, making references to uncomfortable situations and questioning his own actions.
3. Jingjing: Jingjing can be open and candid about his thoughts, even if they seem inappropriate or unusual. He’s also humorously self-aware, undercutting his moments of honesty with reminders that he might be oversharing.
4. Phoebe: Quirky, eccentric, and a free spirit. Her train of thought tends to lean toward the unusual and bizarre. However, she is also compassionate and caring.
4. Yunsheng: Appears offbeat and unusual, suggesting the eating chalk anecdote about his past relationship. He also believes in new-age concepts like auras, showing a more spiritual side.
5. Ross: Insecure and somewhat neurotic and vulnerable. His behavior is indicative of someone who is going through emotional turmoil.
5. Meilin: Exhibits vulnerability and emotional turmoil, especially regarding his recent divorce. He seems to be fluctuating between hurt, anger, and longing for his past relationship.
6. Rachel: Spontaneous and open to change, she takes risks and is adaptable. She is also reliant on her relationships with others, signifying her dependency and need for support.
6. Yusong: Yusong presents as impulsive, high-strung and somewhat comical in moments of panic. Fleeing her wedding because of a sudden realization shows she can make drastic decisions based on her emotions.
E.2 Same-Language Replacement
With Gender-Matched replacement, we replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
in the conversation with
[Sally, Andrew, Ethan, Olivia, Benjamin, Lauren].
With Gender Swap as well, we got [Andrew, Sally,Olivia, Ethan, Lauren,Benjamin].
E.3 Anonymized Speaker Replacement
replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
in the conversation with
[Speaker 0, Speaker 1, Speaker 2, Speaker 3, Speaker 4, Speaker 5]. LLM summarized character description
We use the prompt of "you are a psychology research assistant designed to help analysis character personalities according to conversations……..Here ends the conversation.Give me a character description for each main character according to this conversation." " and we get such result from TVshow Friends scene1.
E.4 Control Test
Example Here is a dialogue between 2 people: Sheldon : (now lifts her head) phoebe!! Amy : you know it’s me? Sheldon : for like a half an hour! man, you can lie about sweden! Amy : how can you come here? Sheldon : how could you not tell me you worked here? Amy : i don’t have to tell you everything!P2: We got an eight-point-two. Trust me, you’re happy.
Prompt The following is a dialogue between the same two people. Please identify who they are:
P0 : (singing) ïpan stripan, glupi glabi! änd that is the swedish national anthem! thank you for asking! (looks annoyed) P1 : wow, ikea… what a rich culture. uhm, you know what? i have a friend who is a masseuse. P0 : oh, ja! ja! P1 : yah! she’s… uhm… not very good though… (phoebe looks devastated)
Table 8: An example for the control test, where in the example,the script is from Friends but speakers are annoted as Amy and Sheldon(from The Big Bang Theory).
Appendix F Verbatim avoid Reasoning
Figure 6: In a personality understanding task where the scenario involves ’Joey returned the lost wallet to its rightful owner, Monica,’ the question posed might be, ’According to this action, his personality should be: ?’ The human reasoning process would draw on contextual evidence, concluding that returning lost property demonstrates honesty. In contrast, the LLM might conclude, ’Joey and Monica are from the TV show Friends, and Joey is characterized as an honest character in this show,’ thus relying heavily on memorized information rather than genuine reasoning.
Appendix G Benchmark Datasets
The six selected benchmarks cover a range of character-centric tasks, designed to test different aspects of character understanding(Table 1):
  • TVShow Character Guessing: Based on the TVSHOWGUESS dataset, which contains scripts from five popular TV series: Friends, The Big Bang Theory, The Office, Frasier, and Gilmore Girls Sang et al. (2022c), this is a multi-choice classification task. The dialogue is anonymized using placeholder identifiers (e.g., P1, P2), and the model must predict the correct speaker for each line, testing its ability to identify characters based on context.
  • Screenplay: Using screenplays from nine popular movies, such as Avengers: Endgame (2019), Dead Poets Society (1989), John Wick (2014), and The Shawshank Redemption (1994) Baruah and Narayanan (2023), this is a coreference resolution task. The model is tasked with linking character mentions throughout the text, forming coreference clusters and resolving ambiguities in character references.
  • Personality Understanding: Using the PERSONET dataset, which comprises approximately 32,000 samples from 33 classic literature books Yu et al. (2023b), this task involves predicting the most likely personality trait depicted in a given text snippet, along with its preceding context. Formulated as a multi-choice problem with distractor traits, this task evaluates the model’s ability to infer character personalities from descriptive text.
  • CSI Role Detection: Based on 39 episodes of the TV series CSI Frermann et al. (2018), this sequence labeling task requires the model to identify mentions of the perpetrator in dialogues and scene descriptions, focusing on role detection within crime-related narratives.
  • FriendsQA: Derived from 1,222 scenes from the first four seasons of Friends Yang and Choi (2019), this dataset supports an open-domain question answering task. The model is given a dialogue excerpt and a manually crafted question and must extract the specific span in the dialogue that answers the question, testing its comprehension of dialogue context.
  • SUMMSCREEN: Built from transcript-recap pairs of 91 TV series collected from community sources such as The TV MegaSite, Inc. (TMS) and ForeverDreaming (FD) Chen et al. (2022), this dataset supports a summarization task. The model is required to generate a coherent summary from dialogue utterances and scene descriptions, capturing key plot points while excluding irrelevant details.
Appendix H Gist Prompting Strategies Across Tasks
We design Gist Prompting methods to reduce reliance on memorization and promote inference-based character understanding. Below, we detail the task-specific prompt formulations.
H.1 Task-Specific Gist Prompts
Character Guessing
"Analyze the dialogue and infer the most likely speaker based on their speech patterns, personality traits, and relationships with other characters. Do not rely on direct recall of specific names but rather use reasoning from the provided context."
Coreference Resolution
"Identify which mentions in the text refer to the same character by analyzing linguistic patterns, contextual cues, and discourse coherence. Do not rely on pre-learned entity associations but rather resolve coreference based on narrative structure."
Personality Understanding
"Based on the character’s dialogue and actions, infer their personality traits. Focus on how they interact with others, their tone, and behavioral tendencies. Avoid using prior memorized descriptions of characters and derive insights purely from the given passage."
Role Detection
"Determine whether a character plays a key role in the narrative (e.g., antagonist, victim, or protagonist) by analyzing their function within the story. Identify how their actions and relationships define their role rather than relying on predefined labels."
Open-Domain QA
"Answer the question using only the given passage. Identify relevant information by reasoning over the text and analyzing interactions between characters. Do not rely on external knowledge or prior memorized facts."
Summarization
"Summarize the passage by identifying key events, character interactions, and narrative progression. Construct a coherent summary that captures the essence of the scene without directly copying text segments."