### v5

# Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation

Abstract
Large Language Models (LLMs) achieve striking performance on character understanding tasks, yet this success raises a key concern: are they truly reasoning about fictional characters, or merely recalling them from memorization? While prior work has explored memorization in general reasoning tasks, it remains unclear how memorization affects character understanding. To investigate this, we draw on cognitive science distinctions between "verbatim" (exact recall) and "gist" (semantic abstraction) memorization, and introduce a graded intervention framework with two levels of disruption. The soft setting preserves surface text but applies prompting strategies to encourage reasoning over recall, while the hard setting perturbs key character references to directly block memorization cues. Evaluating GPT-4o, LLaMA3.3-70B, and DeepSeek V3 on six character understanding benchmarks, we find that our hard setting leads to a dramatic performance drop of 45.2%—revealing substantial reliance on verbatim memorization and data leakage in existing benchmarks. In contrast, the soft setting steers models to reason more robustly without harming performance, showing that prompting alone can shift LLMs between memorization-based and reasoning-based strategies.
Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation
Yuxuan Jiang , Francis Ferraro Department of Computer Science and Electrical Engineering University of Maryland, Baltimore County yuxuanj1@umbc.edu
1 Introduction
Character understanding tasks assess a model’s ability to comprehend fictional characters’ personalities, goals, roles, and relationships within a narrative Yu et al. (2023a); Vallurupalli et al. (2024); Shen et al. (2024); Brahman et al. (2021). Recent LLMs have shown strong performance on such benchmarks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023), raising an important question: do these models truly understand characters, or are they simply recalling text seen during pretraining Chang et al. (2023a)? For instance, in a speaker guessing task from The Big Bang Theory (Figure 1), a model may identify Sheldon correctly not by reasoning over past actions (e.g., preparing lessons for Howard), but by matching exact phrases seen in previous data.
This concern is especially relevant in character understanding, as many benchmarks are based on widely consumed fictional works, such as TV shows, movies, or novels Sang et al. (2022a). Given the massive and opaque pretraining corpora of LLMs, models are likely to have encountered these works verbatim. As illustrated in Figure 1, such memorization can yield high performance even without genuine understanding. As a result, performance may reflect reliance on what cognitive psychology calls verbatim memorization—exact recall of surface text Huang et al. (2024); Ippolito et al. (2022)—rather than gist memorization, which captures semantic understanding and general patterns Reyna and Brainerd (1998). This challenges the validity of current evaluations: are models demonstrating true reasoning, or simply relying on verbatim memorization to regurgitate answers? In this work, we formulate a testable hypothesis: if LLMs’ strong performance is largely driven by memorized surface cues, then removing those cues—such as character names—should lead to a significant drop in accuracy.
Figure 1: Gist vs. verbatim memorization in speaker identification under original and anonymized scripts. This figure compares two memorizaiton strategies in identifying speakers (P0, P1) from a dialogue. On the left, character names are intact; on the right, all names are anonymized. The gist-based approach (blue) is prompted to "analyze personality traits, tones, and relationships", and infers speaker identity through reasoning—remaining robust even when names are masked. In contrast, the verbatim-based approach (yellow) is prompted to "recall names from exact dialogue memorization", and may misattribute the script (e.g., predicting Friends instead of The Big Bang Theory) when superficial cues are removed.
To investigate this hypothesis, we develop a graded intervention framework with two levels of disruption: a soft setting and a hard setting. The soft setting retains the original input scripts but uses gist-based prompting to guide the model toward abstraction and inference, without removing surface-level cues. In contrast, the hard setting replaces all character names with culturally distinct alternatives, aiming to block verbatim memorization while preserving the narrative content. We apply this graded intervention framework to six character understanding tasks and evaluate three LLMs: GPT-4o, LLaMA3-70B, and DeepSeek V3.
Our experiments provide clear support for our initial hypothesis: removing verbatim memorization cues significantly reduces LLMs’ performance on character understanding tasks, confirming that models heavily rely on memorized surface content. Specifically, we observe performance drops of up to 45.8% when verbatim memorization cues are blocked. In contrast, gist-based prompting maintains accuracy by steering model behavior toward reasoning rather than verbatim recall. These results suggest that current benchmarks may overestimate models’ genuine reasoning capabilities and that prompting strategies can effectively disentangle reasoning from memorization.
In summary, our key contributions are:
  1. 1.
    Identifying Memorization in Character Understanding: We empirically demonstrate significant performance drops (up to 46.2%) across benchmarks when verbatim memorization cues are disrupted, revealing a heavy reliance on surface-level recall rather than genuine reasoning.
  2. 2.
    A Graded Framework to Disentangle Reasoning from Memorization: Inspired by cognitive distinctions, we introduce a soft–hard intervention framework. The soft setting employs gist-based prompting to encourage abstraction, while the hard setting explicitly blocks verbatim recall through surface-level perturbation (Section 6).
  3. 3.
    Prompting as a Diagnostic Tool: We show that prompting style profoundly influences model behavior: verbatim prompts reinforce memorization, whereas gist prompts facilitate deeper inference, providing a practical approach to probe underlying reasoning processes in LLMs (Section 4.3).
2 Related Work
2.1 Character Understanding and Its Evaluation with LLMs
Character understanding has long been a focus in psychological, literary, and educational research Bower and Morrow (1990); McKee (2010); Paris and Paris (2003). It spans tasks such as character recognition, coreference resolution, summarization, role detection, goal modeling, personality analysis, and question answering Bamman et al. (2020); Chen et al. (2022); Frermann et al. (2018); Vallurupalli et al. (2024); Yu et al. (2023b).
Recent work demonstrates that LLMs achieve strong results on character-centric benchmarks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023). However, many of these datasets are derived from publicly available or widely consumed fictional sources, such as movie scripts, TV shows, and novels Sang et al. (2022a); Tapaswi et al. (2016); Caselli and Vossen (2017), which makes them vulnerable to pretraining overlap and surface-level memorization Chang et al. (2023a). Although prior research has studied LLM memorization broadly Carlini et al. (2023); Xie et al. (2024); Yang et al. (2023a), little is known about how it specifically affects character understanding tasks. Our work fills this gap by examining memorization’s role in these tasks and proposing interventions to distinguish between recall and reasoning.
2.2 Memorization and Reasoning in LLMs
LLMs’ impressive reasoning capabilities have attracted significant attention Tong et al. (2023); Yang et al. (2023b); Xiong et al. (2024); Tan et al. (2024); Jiang et al. (2025), yet their behavior is also strongly influenced by memorization Zhou et al. (2023); Chen et al. (2024); Biderman et al. (2024). Prior studies indicate that memorization of training data can artificially inflate benchmark results Khandelwal et al. (2020); Blevins and Zettlemoyer (2022), while conflicting memorized content can negatively affect performance Su et al. (2024); Wang et al. (2024); Yang et al. (2025). Unlike structured reasoning tasks, character understanding tasks inherently involve some memorization (e.g., narrative consistency), complicating the distinction between genuine reasoning and mere recall.
Cognitive science provides a useful conceptual framework for analyzing memory, distinguishing between verbatim memorization—the precise recall of surface-level details—and gist memorization—the abstraction of semantic meaning Reyna and Brainerd (1998); Brainerd and Reyna (2002). To avoid confusion with architectural notions of "memory" in NLP and machine learning Alizadeh et al. (2024); Kwon et al. (2023), we consistently adopt these terms to describe different recall styles rather than neural memory mechanisms. In our work, this distinction serves as the theoretical foundation for disentangling shallow memorization from deeper reasoning behavior in large language models.
Recent NLP research has formalized methods for quantifying and disentangling memorization from reasoning Lesci et al. (2024); Jin et al. (2024); Carlini et al. (2023); Prabhakar et al. (2024), examining phenomena such as multifaceted memorization Prashanth et al. (2024); Li et al. (2024) and logit-based tracing Chen et al. (2024). Mitigation strategies (e.g., out-of-distribution filtering Biderman et al. (2024)) are challenging to apply to character-centric tasks, which typically rely on known fictional content. Thus, we propose a prompting-based approach designed specifically to reduce verbatim reliance and promote gist-level reasoning in character understanding tasks.
3 Data and Experiment Setup
To comprehensively evaluate character understanding in LLMs, we compile two datasets: CharBench and CharScript. Inspired by NLP benchmark collections such as GLUE Wang et al. (2018), our CharBench integrates six publicly available datasets spanning character-centric tasks like coreference resolution, personality inference, and dialogue-based QA (details in Section 6.1).
The second dataset, CharScript, is newly constructed for this study by systematically sampling textual segments from original fictional sources underlying CharBench, explicitly probing models’ verbatim memorization (details in Section 3.2).
3.1 CharBench: A Benchmark Collection for Character Understanding
CharBench aggregates six widely recognized NLP benchmarks covering diverse narrative mediums—TV shows, movies, and novels. To ensure comparability, we retain each dataset’s original splits, preprocessing, and evaluation metrics. Detailed task descriptions and implementations are provided in Appendix E.
Benchmark Sources
Following established NLP benchmark practices Wang et al. (2018), CharBench includes:
  • MovieCoref Chen and Choi (2016): Coreference resolution in movie scripts.
  • TVSHOWGUESS Sang et al. (2022b): Speaker identification in TV dialogues.
  • PERSONET Yu et al. (2023b): Personality inference from literary passages.
  • CSI Corpus Frermann et al. (2018): Role detection from crime shows.
  • FriendsQA Yang and Choi (2019): Multi-turn dialogue-based QA.
  • SummScreen Chen et al. (2022): TV episode summarization.
Evaluation Metrics
We strictly follow original dataset metrics (exact match, ROUGE, F1 score), as detailed in Table 1.
Task Name Task Type Data Source Metric
TVShow Guess Character Guessing TV Shows Exact Match
FriendsQA Question Answering TV Shows Exact Match
Screenplay Character Coreference Movies F1 Score
SummScreen Text Summarization TV Shows ROUGE
CSI Role Detection TV Shows F1 Score
PERSONET Personality Understanding Novels Exact Match
Table 1: Overview of benchmark tasks, including their type, data sources, and evaluation metrics.
3.2 CharScript: Fictional Works Memorization Benchmark Dataset
Motivated by recent NLP research on memorization in language models Chang et al. (2023b), we constructed the CharScript dataset to evaluate whether LLMs can recognize the source of fictional text segments. Following a similar setup to Chang et al. (2023b), we probe whether models can correctly attribute short passages to their original narrative sources, as an indicator of verbatim memorization. We present our experimental design and findings in Section 5.1.
We systematically sampled and extracted 200 textual segments, averaging approximately 312 words each, from 40 diverse fictional sources (11 TV series, 9 movies, and 20 novels). These sources were carefully selected to maximize narrative diversity, textual complexity, and representativeness of popular fictional works included in CharBench. Each source contributes exactly 5 segments, structured uniformly in the format [Title]; Scene context, combining both original dialogue and narrative descriptions.
3.3 Models
We evaluate three frontier LLMs—open-source LLaMA 3.3-70B AI (2024) and DeepSeek V3 Liu et al. (2024), as well as the closed-source GPT-4o OpenAI (2023)—via API-based inference. To address stochastic variability, each model was run three times per setting, and average performance is reported. The total cost of our experiments was approximately \$100.
4 Soft Setting: Memorization-Controlled Prompting in Character Guessing
In the introduction, we formulated a testable hypothesis that LLMs’ strong performance on character understanding tasks relies significantly on verbatim memorization. As a first step toward testing this hypothesis, we investigate whether prompt design alone—without explicitly removing surface-level textual cues—can shift model responses from verbatim recall toward gist-based reasoning.
4.1 Character Guessing Task Format
We select the Character Guessing Task specifically because it closely mirrors the name cloze tests Chang et al. (2023b) previously established as effective probes for verbatim memorization in LLMs. Name cloze tests measure memorization by removing character names from familiar contexts and testing if models can accurately recall them. Similarly, in our task, speaker labels are removed from dialogue excerpts, requiring the model to infer speakers solely from textual context. If models rely heavily on verbatim memorization, accuracy should significantly degrade when surface cues—such as explicit character names—are disrupted. Conversely, if models genuinely reason about character interactions and relationships, performance should remain relatively stable despite these perturbations.
Following the original TVShow Guess setup Sang et al. (2022c), the model is tasked with predicting the speaker of each utterance in a dialogue segment where all speaker labels are removed.
4.2 Prompting Strategies
We design two distinct prompting conditions to examine how LLMs rely on different memorization mechanisms when identifying characters. Verbatim prompting encourages mechanical memorization, where the model recalls exact substrings from its training data and reconstructs missing parts without reasoning. In contrast, gist prompting focuses on extracting key information and performing memorization-based matching. For character understanding tasks, the most relevant gist cues typically include character relationships, significant events, and personality traits.
All prompts share the same base instruction: "The following is a dialogue. Please identify who they are. …"
The two variations are as follows:
Verbatim Prompt: "…Do this by direct memorization retrieval—do not infer based on traits, interactions, or context. Recall names exactly from past dialogue knowledge without explanations or reasoning."
Gist Prompt: "…Do this by analyzing relationships, key events, and personality traits to match them with known characters. Do not directly retrieve from memorization who said the sentences."
4.3 Result and Analysis
Our key observations are as follows:
  • Baseline Performance: The baseline condition reflects the overall performance of the models, with GPT-4o demonstrating a clear lead to others, while Llama 3.3-70B and DeepSeek V3 exhibit quite similar performance.
  • Memorization Reliance: All three models show that the Verbatim condition performs nearly identically to the baseline, supporting our hypothesis that models primarily rely on memorization when answering the Character Guessing task.
  • Effect of Gist Prompting: The Gist condition results in a non-trivial performance drop, approximately 10%, suggesting that our prompt effectively reduces the model’s dependence on memorization when generating responses.
Our results indicate that gist-based prompting can indeed shift model responses toward reasoning, as evidenced by the performance drop under the Gist condition (see Figure 2). However, the moderate performance drop (approximately 10%) suggests that even gist prompting may not completely eliminate the models’ reliance on verbatim memorization. Previous research similarly indicates that prompting without explicit reasoning steps may fail to fully engage true reasoning processes Yu (2025). Therefore, to conclusively test our hypothesis—that models’ strong benchmark performance relies significantly on verbatim memorization—we introduce a more direct and explicit intervention aimed at disrupting memorized surface-level cues.
Figure 2: The bar chart compares the performance of three LLMs under three prompting settings. Across all three models, the performance of the Verbatim and Baseline conditions is highly similar, confirming our hypothesis that the models rely on memorization when answering questions. The Gist condition, however, shows a non-trivial decline in performance, indicating that the models struggle when they cannot fully rely on memorization to generate responses.
5 Hard Setting: Memorization-Controlled Prompting under Name Replacement
To effectively suppress verbatim memorization and induce gist-style reasoning, we design a stronger intervention: replacing character names to explicitly remove memorized surface cues. While prior work uses name replacement to improve model robustness Yu et al. (2022), we take this idea further by systematically testing whether such name-based interventions are strong enough to disrupt verbatim memorization and induce gist-based reasoning.
5.1 Name Replacement Strategies and Their Impact on Memorization
Name Replacement Strategies
Previous studies Yu et al. (2022) have primarily employed Same-Cultural Name Replacement, a naive and intuitive approach that substitutes character names with culturally consistent alternatives. To systematically analyze the impact of name replacement on model performance, we introduce three distinct levels of name modifications:
  • Name Masking: Replacing character names with placeholders (e.g., Sheldon, Amy → P0, P1).
  • Cross-Cultural Name Replacement: Substituting names with those from a different cultural background (e.g.English → Chinese) while maintaining gender consistency (e.g., Sheldon, AmyJie Zhang, Xiaoling WangSun et al. (2024). This tests whether models rely on specific name distributions for character inference.
  • Same-Cultural Name Replacement: Replacing names with other culturally consistent alternatives (e.g., Sheldon, AmyAndrew, Sally). This retains cultural familiarity while disrupting direct memorization by name replacements.
Memorization Evaluation Task Format
To directly quantify how much verbatim memorization models rely on, we design a Source Identification Task. This task closely resembles established memorization evaluation methods, such as source attribution or textual cloze tests Chang et al. (2023b). Specifically, given a dialogue excerpt $D=\{U_{1},U_{2},...,U_{n}\}$ without source annotations, the model must identify the correct origin of the text, predicting $f_{\text{source}}(D)\rightarrow S$, where $S$ is the specific fictional source.
The rationale behind this design is straightforward: if a model has memorized verbatim text from pretraining, it should accurately identify the source based purely on surface-level textual recall. Conversely, if name replacement or prompting disrupts memorization effectively, we expect a significant decrease in source-identification accuracy. Thus, observing substantial performance degradation would strongly support the claim that models’ prior accuracy was driven largely by verbatim memorization rather than reasoning. We use the CharScript: Fictional Works Memorization Benchmark Dataset described in Section 3.2 to rigorously evaluate this effect.
Impact of Name Replacement on Memorization.
According to Figure 3, our results demonstrate that different name replacement strategies have a significant impact on the memorization ability of LLMs. Among them, Cross-Cultural Name Replacement consistently induces the strongest forgetting effect across all models, indicating that altering names to those from a different cultural background disrupts memorized associations more effectively than other replacement methods. This suggests that cultural distance may play a crucial role in weakening memorized identity-linking patterns in models.
Figure 3: The heatmap presents accuracy percentages for source identification across three LLMs under different name replacement conditions. All three models demonstrate strong memorization of the script. The Cross-Culture Name Replacement method exhibits a significantly strong and the most pronounced memorization-weakening effect across all three models.
5.2 Memorization-Controlled Prompting under Name Replacement
Different name replacement strategies have shown a significant impact on the memorization ability of LLMs. Building on this, we apply the name replacement strategies introduced in Section 5.1 to evaluate the Character Guessing tasks, using the verbatim and gist prompting strategies we have discussed in Section 4.2.
Under the Name Replacement condition, we provide brief character descriptions generated by GPT-4, aiming to enable the model to complete the Character Guessing task through character matching. The descriptions used are provided in the Appendix C. To assess the actual effectiveness of these descriptions, we conduct an ablation study.
Figure 4: The line chart shows the decline in memorization across different settings, illustrating that as memorization weakens, overall model performance also decreases. The bar chart presents the effect of different prompting strategies: under Name Replacement conditions, Verbatim prompting consistently outperforms the baseline by reinforcing memorization, whereas Gist prompting further reduces reliance on memorization, leading to the lowest accuracy.
Results and Key Findings
We present the core results in Figure 4 and summarize four key findings:
  1. 1.
    Performance Declines with Reduced Memorization. Accuracy drops as surface cues are removed, revealing the model’s strong reliance on verbatim memorization.
  2. 2.
    Character Descriptions Have Minimal Impact. Adding character descriptions does not notably change performance, suggesting limited influence on model decisions.
  3. 3.
    Prompting Style Matters. Verbatim prompts boost performance by reinforcing memorization, while Gist prompts suppress it and lead to lower accuracy.
  4. 4.
    Name Replacement + Gist Prompting Is Effective. This combination best reduces memorization reliance and reveals the model’s true reasoning ability.
6 Evaluating Character Understanding Tasks Under Memorization Constraints
Sections 4 and 5 provided clear empirical evidence supporting our hypothesis: LLMs’ strong performance on character understanding tasks significantly relies on verbatim memorization. To further examine the generalizability of our approach and assess how memorization affects performance across diverse task types, we now extend the combination of Gist Prompting and Cross-Cultural Name Replacement to five additional character understanding benchmarks.
6.1 Benchmark Tasks Description
We benchmark five types of character understanding tasks spanning coreference, personality inference, role detection, QA, and summarization. Each task evaluates the model’s ability to extract or infer character-related information from dialogue.
Example Dialogue
To illustrate these tasks, consider the following conversation:
P0: "You’re still going to have to convince a jury that I killed two strangers for no reason."
Grissom doesn’t look worried. He takes his gloves off and puts them on the table.
P1: "You ever been to the theater, Peter? There’s a play called Six Degrees of Separation."
The correct speaker labels are P0 → Peter Berglund, P1 → Grissom. For all tasks except Character Guess, the placeholders P0 and P1 are replaced with their original names, Peter Berglund and Grissom. Table 2 illustrates how this dialogue segment is utilized across different tasks.
Task Example Question and Expected Output
Coreference Resolution "Which entity does ’his’ refer to?" → "Grissom"
Personality Understanding "What trait best describes P1?" → Analytical
Role Detection "Identify all entities referring to the criminal." → Peter Berglund
Question Answering "What play does Grissom mention?" → "Six Degrees of Separation"
Summarization "Generate a summary." → "Peter discusses convincing a jury, while Grissom references a play."
Table 2: Examples of different character understanding tasks applied to the same dialogue.
6.2 Gist Prompts Across Character Understanding Tasks
General Guidelines for Gist Prompting
  • Encourage Structural Reasoning: Guide models to infer character identity and relationships from discourse patterns rather than relying on explicit name associations.
  • Eliminate Memorization Bias: Ensure prompts emphasize evidence-based inference rather than knowledge retrieval.
  • Preserve Narrative Coherence: In summarization and question answering, focus on maintaining logical flow rather than phrase-matching.
Detailed prompts in Appendix F.
{NiceTabular}
l|ccc|ccc|ccc Tasks GPT-4o LLaMA3.3-70B DeepSeek V3
Origin NR NR+GIST Origin NR NR+GIST Origin NR NR+GIST
Tvshow Guess 78.2 32.0 $\downarrow$46.2 31.5 $\downarrow$0.5 61.4 31.9 $\downarrow$29.5 27.8 $\downarrow$4.1 60.8 29.7 $\downarrow$31.1 26.7 $\downarrow$3.0
Coreference 58.3 49.6 $\downarrow$8.7 47.5 $\downarrow$2.1 56.7 48.3 $\downarrow$8.4 48.1 $\downarrow$0.2 47.4 34.6 $\downarrow$12.8 32.3 $\downarrow$2.3
FriendsQA 44.3 38.6 $\downarrow$5.7 38.6 $\uparrow$0.0 45.7 42.9 $\downarrow$2.8 44.3 $\uparrow$1.4 42.8 39.4 $\downarrow$3.4 34.3 $\downarrow$5.1
ScreenSum 37.5 28.7 $\downarrow$8.8 23.1 $\downarrow$5.6 40.1 25.9 $\downarrow$14.2 21.3 $\downarrow$4.6 35.6 19.3 $\downarrow$16.3 17.1 $\downarrow$2.2
CSI Role Extract 50.1 32.9 $\downarrow$17.2 31.7 $\downarrow$1.2 48.3 34.7 $\downarrow$13.6 32.7 $\downarrow$2.0 45.1 29.8 $\downarrow$15.3 23.7 $\downarrow$6.1
PERSONET 56.0 52.0 $\downarrow$4.0 54.0 $\uparrow$2.0 62.0 46.0 $\downarrow$16.0 44.0 $\downarrow$2.0 44.0 40.0 $\downarrow$4.0 38.0 $\downarrow$2.0
Table 3: Ablation study examining the impact of name replacement and the GIST method on model performance across six NLP tasks.The subscripted values in the NR column indicate the performance drop relative to the original script, while the subscripted values in the NR+GIST column show the performance change compared to the Name Replace condition.
6.3 Results and Key Findings
Our evaluation reinforces our conclusions regarding memorization versus reasoning, with three key findings:
  1. 1.
    Significant Performance Drop with Reduced Memorization: As shown in Table 6.2, applying the name replacement (NR) method leads to a substantial performance decline across all tasks. This confirms our hypothesis that current character understanding benchmarks are heavily reliant on explicit memorization, potentially inflating standard evaluation scores. Furthermore, prompting models to use gist reasoning induces an additional performance drop, underscoring the effectiveness of gist prompting in mitigating reliance on surface-level memorization.
  2. 2.
    Performance Gaps across Task Types (Motivation vs. Action Prediction): We observe clear differences in memorization effects between action-oriented tasks (e.g., Character Guessing, CSI Role Detection) and motivation-oriented tasks (e.g., Personality Understanding). Action prediction tasks typically require models to output discrete labels—such as a character name or role—making it easy for models to memorize associations between surface cues and labels. For instance, phrases consistently linked to characters like Sheldon (“I’m not crazy, my mother had me tested”) are easily memorized, but become brittle under name replacement. In contrast, motivation prediction tasks involve deeper contextual inference about characters’ internal states and intentions, rendering surface-level memorization less effective, thus promoting genuine reasoning.
  3. 3.
    Memorization Can Be Detrimental: Interestingly, certain tasks show performance improvements under the NR+Gist setting. Our error analysis reveals that reducing memorization resolves knowledge conflicts—cases where memorized but incorrect associations previously misled models, indicating that excessive memorization can actively impair true character understanding.
These findings are consistent across multiple models (Figure 5), suggesting generalizability and broader implications of memorization biases in current benchmarks.
Figure 5: Performance degradation across LLMs due to GIST+NR. Action prediction tasks (e.g., Guessing, CSI) show greater drops than motivation prediction tasks (e.g., PERSONET, FriendsQA), reflecting their differential reliance on memorization.
7 Conclusion
In this work, we introduce the distinction between verbatim memorization and gist memorization, and systematically demonstrate that many existing character understanding evaluations rely heavily on verbatim memorization, potentially inflating perceived reasoning capabilities of LLMs. To rigorously test and mitigate this reliance, we propose a novel, graded intervention framework combining Gist Prompting and Cross-Cultural Name Replacement. Our experimental results across six diverse character-centric benchmarks clearly show that reducing verbatim memorization cues significantly affects model performance, highlighting the importance of explicitly disentangling memorization from genuine reasoning.
Our findings emphasize the critical need for developing memorization-aware evaluations, particularly in domains like character understanding where memorization effects are pronounced. Broadly, our methodology offers generalizable insights applicable to a wide range of inference tasks where accurately distinguishing between memorization and reasoning is essential.
8 Limitations
The specific training data of GPT-4o is not publicly disclosed by OpenAI, making it fundamentally unknowable. Our work employs probabilistic inference to estimate the model’s familiarity with a set of fictional works. However, whether these works are explicitly included in the training data remains unanswerable. Nevertheless, even mere familiarity can suffice to contaminate test data, posing challenges to fair evaluation.
Our study can be extended by future research in a number of ways. First, though we ground and motivate our work through “verbatim” and “gist” memorization, we note that any apparent “reasoning” of AI systems may differ fundamentally from those of humans, and we did not pursue human studies to examine this in detail. Further investigation is needed to explore this topic.
Second, the relationship between memorization and reasoning remains a subject of debate in cognitive science. Future studies may uncover insights into this relationship from entirely new perspectives.
Thirdly, While our interventions reveal strong trends, we acknowledge that results may be sensitive to the surface form of prompts. Future work could complement our findings with uncertainty-based analyses such as log probability comparisons.
Finally, our study was centered on select, recent (but not contemporaneous) majority Western popular culture references. Given their popularity, this selection likely increases the chances that LLMs have been pretrained on more instances or language discussing those references (vs. contemporaneous references, or references from non-Western cultures, where overall occurrences in training data may be lower); it is possible that this exaggerated some of the effects we observed.
9 Ethics
Our study probes the degree to which GPT-4o have memorized fictional works, and its impact on character understanding tasks. This work uses the OpenAI API on experiments above, and at no point do we access, or attempt to access, the true training data behind these models, or any underlying components of the systems. Risks The several datasets and fictional work scripts in our experiment are sourced from publicly available sources. However, we cannot guarantee that they are free from socially harmful or toxic language. We use ChatGPT  [^1] https://chatgpt.com/ to correct grammatical errors in this paper.
Appendix A Example of Script-based Character Understanding Task
In Table˜4 we show a full example of the character understanding task. In this example, we have automatically replaced the individual character names with placeholders like P0 and P1.
Input P0: Hey, sorry about that P1: No, we’re sorry. We never should have been comparing relationships in the first place. P2: Why? We won. You know, I say, next, we take on Koothrappali and his dog. Really give ourselves a challenge. P3: I just want to say one more thing about this. Just because Penny and I are very different people does not mean that we’re a bad couple. P2: The answer is one simple test away. Hmm? You know, it’s like when I thought there was a possum in my closet. Did I sit around wondering? No, I sent Leonard in with a pointy stick and a bag. P3: I killed his Chewbacca slippers. P0: Let’s just take the test. P3: No, no, no, I don’t want to. P0: Oh, well, ’cause you know we’re gonna do bad. P3: Because it doesn’t matter. I don’t care if we’re a ten or a two. P2: Or a one. A one is possible. P3: Marriage is scary. You’re scared, I’m scared. But it doesn’t make me not want to do it. It, it just makes me want to hold your hand and do it with you. P0: Leonard. P1: It makes me so happy if you said things like that. P2: We got an eight-point-two. Trust me, you’re happy.
Label P0: Penny, P1: Amy, P2: Sheldon, P3: Leonard
Table 4: An example from character guessing task.
Input Tell me the source for following script:Penny: Hey, sorry about that Amy: No, we’re sorry. We never should have been comparing relationships in the first place. Sheldon: Why? We won. You know, I say, next, we take on Koothrappali and his dog. Really give ourselves a challenge. Leonard: I just want to say one more thing about this. Just because Penny and I are very different people does not mean that we’re a bad couple.
Source The Big Bang Theory
Table 5: An example for source prediction task.
Appendix B Extended Dataset Analysis
For the six representative character understanding tasks studied in our work, we selected one dataset for each category. However, we also recognize that there are many different benchmark datasets for each type of task. When selecting representative datasets for each category, we focused on choosing those with more diverse sources. Additionally, we supplemented the selection with eight datasets of the same type but derived from different sources in Table 6.
Task Name Task Type Source
Character Mining Character Identification TV Shows
LISCU Character Identification Book Reviews
MovieQA Visual Question Answering Movies
Liter Coreference Character Coreference Novels
HVV Detect Role Detection Movies, News
SYMON Visual Question Answering Movies
CEB Character Representation Novels
Love and Violence Summarization TV Shows
Table 6: Overview of 8 more tasks in the character understanding benchmark.
Appendix C Prompts for GPT4
The following is the prompt we sent to GPT4: For personality traits extraction ’you are a psychology research assistant designed to help analysis character personalities according to conversations.’ conversation ’Here ends the conversation.Give me a character description for each main character according to this conversation.’
Character Description Here is a character description:Monica: Responsible, caring, and organized. Assertive and confident in her actions.Chandler: Witty and self-deprecating with an approachable sense of humor. Exhibits insecurity and anxiety in his dialogues, making references to uncomfortable situations and questioning his own actions.Rachel: Spontaneous and open to change, she takes risks and is adaptable. She is also reliant on her relationships with others, signifying her dependency and need for support.’ ’Joey: Energetic, extroverted, and casual. Lacks the sensitivity of others feelings at times but genuinely care about friends.Phoebe: Quirky, eccentric, and a free spirit. Her train of thought tends to lean toward the unusual and bizarre. However, she is also compassionate and caring.Ross: Insecure and somewhat neurotic and vulnerable. His behavior is indicative of someone who is going through emotional turmoil.’
Appendix D Name Replacement and Character Comparison
We also provide partial name replacement dictionary in table  7.
D.1 Cross-Language Uncommon Name Replacement
With gender swap:[Bojing, Cuixia, Jingjing, Yunsheng, Meilin, Yusong] Without gender swap:[Cuixia, Bojing,Yunsheng,Jingjing,Yusong, Meilin]
Task name Before replacement After replacement
TVShow Monica, Joey, Chandler, Phoebe, Ross, Rachel Bojing, Cuixia, Jingjing, Yunsheng, Meilin, Yusong
CSI Grissom, Catherine, Nick, Sara, Brass Jingbo, Jingwen, Huaqiang, Lihui, Weimin
PERSONET Villefort, The Count, Dantès, Caderousse, Danglars, Albert, Mercédès, Elders, Abbé Busoni, Eugénie, Faria priest, Fernand, Franz, Faria, Morre, Monte Cristo Weina, Nüxia, Danqing, Cailing, Danyu, Ailan, Mengzhe, Lianhua, Busu, Yujun, Fanli, Feiyan, Fangqing, Fangli, Moya, Mingxia
Table 7: Besides the main characters, there are many names of minor characters and passersby appearing in the tasks. In each task, these secondary and passerby names can be dozens of times more than the number of main characters. Replacing these names is complex but not impossible (e.g., with the help of LLMs). However, to align with the TV show approach of only replacing the names of main characters, we only replaced the main characters in all tasks. From the results, it is clear that even replacing only the main characters has had a significant effect.
replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
in the conversation with
[Bojing, Cuixia, Jingjing, Yunsheng, Meilin, Yusong].
The comparison before and after name substitution
The scene overall reflects a group dynamic characteristic of close friends who are comfortable jesting with each other, while also being there in times of crisis or emotional turmoil.
1.Monica: Responsible, caring, and organized. Assertive and confident in her actions.
1. Bojing: Bojing comes across as practical, level-headed and caring. He often acts as the voice of reason for his friends, attempting to mediate, clarify, and console in various situations. His attempts to play down his date suggest he is a private person who doesn’t enjoy sharing intimate details of his life.
2. Joey: Energetic, extroverted, and casual. Lacks the sensitivity of others’ feelings at times but genuinely care about friends.
2. Cuixia: Cuixia seems like a lively, fun-loving character. However, she sometimes shows a more cynical side, quick to suspect something might be wrong with Bojing’s date and suggesting a strip joint as a solution to Meilin’s woes.
3. Chandler: Witty and self-deprecating with an approachable sense of humor. Exhibits insecurity and anxiety in his dialogues, making references to uncomfortable situations and questioning his own actions.
3. Jingjing: Jingjing can be open and candid about his thoughts, even if they seem inappropriate or unusual. He’s also humorously self-aware, undercutting his moments of honesty with reminders that he might be oversharing.
4. Phoebe: Quirky, eccentric, and a free spirit. Her train of thought tends to lean toward the unusual and bizarre. However, she is also compassionate and caring.
4. Yunsheng: Appears offbeat and unusual, suggesting the eating chalk anecdote about his past relationship. He also believes in new-age concepts like auras, showing a more spiritual side.
5. Ross: Insecure and somewhat neurotic and vulnerable. His behavior is indicative of someone who is going through emotional turmoil.
5. Meilin: Exhibits vulnerability and emotional turmoil, especially regarding his recent divorce. He seems to be fluctuating between hurt, anger, and longing for his past relationship.
6. Rachel: Spontaneous and open to change, she takes risks and is adaptable. She is also reliant on her relationships with others, signifying her dependency and need for support.
6. Yusong: Yusong presents as impulsive, high-strung and somewhat comical in moments of panic. Fleeing her wedding because of a sudden realization shows she can make drastic decisions based on her emotions.
D.2 Same-Language Replacement
With Gender-Matched replacement, we replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
in the conversation with
[Sally, Andrew, Ethan, Olivia, Benjamin, Lauren].
With Gender Swap as well, we got [Andrew, Sally,Olivia, Ethan, Lauren,Benjamin].
D.3 Anonymized Speaker Replacement
replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
in the conversation with
[Speaker 0, Speaker 1, Speaker 2, Speaker 3, Speaker 4, Speaker 5]. LLM summarized character description
We use the prompt of "you are a psychology research assistant designed to help analysis character personalities according to conversations……..Here ends the conversation.Give me a character description for each main character according to this conversation." " and we get such result from TVshow Friends scene1.
Appendix E Benchmark Datasets
The six selected benchmarks cover a range of character-centric tasks, designed to test different aspects of character understanding(Table 1):
  • TVShow Character Guessing: Based on the TVSHOWGUESS dataset, which contains scripts from five popular TV series: Friends, The Big Bang Theory, The Office, Frasier, and Gilmore Girls Sang et al. (2022b), this is a multi-choice classification task. The dialogue is anonymized using placeholder identifiers (e.g., P1, P2), and the model must predict the correct speaker for each line, testing its ability to identify characters based on context.
  • Screenplay: Using screenplays from nine popular movies, such as Avengers: Endgame (2019), Dead Poets Society (1989), John Wick (2014), and The Shawshank Redemption (1994) Baruah and Narayanan (2023), this is a coreference resolution task. The model is tasked with linking character mentions throughout the text, forming coreference clusters and resolving ambiguities in character references.
  • Personality Understanding: Using the PERSONET dataset, which comprises approximately 32,000 samples from 33 classic literature books Yu et al. (2023b), this task involves predicting the most likely personality trait depicted in a given text snippet, along with its preceding context. Formulated as a multi-choice problem with distractor traits, this task evaluates the model’s ability to infer character personalities from descriptive text.
  • CSI Role Detection: Based on 39 episodes of the TV series CSI Frermann et al. (2018), this sequence labeling task requires the model to identify mentions of the perpetrator in dialogues and scene descriptions, focusing on role detection within crime-related narratives.
  • FriendsQA: Derived from 1,222 scenes from the first four seasons of Friends Yang and Choi (2019), this dataset supports an open-domain question answering task. The model is given a dialogue excerpt and a manually crafted question and must extract the specific span in the dialogue that answers the question, testing its comprehension of dialogue context.
  • SUMMSCREEN: Built from transcript-recap pairs of 91 TV series collected from community sources such as The TV MegaSite, Inc. (TMS) and ForeverDreaming (FD) Chen et al. (2022), this dataset supports a summarization task. The model is required to generate a coherent summary from dialogue utterances and scene descriptions, capturing key plot points while excluding irrelevant details.
Appendix F Gist Prompting Strategies Across Tasks
We design Gist Prompting methods to reduce reliance on memorization and promote inference-based character understanding. Below, we detail the task-specific prompt formulations.
F.1 Task-Specific Gist Prompts
Character Guessing
"Analyze the dialogue and infer the most likely speaker based on their speech patterns, personality traits, and relationships with other characters. Do not rely on direct recall of specific names but rather use reasoning from the provided context."
Coreference Resolution
"Identify which mentions in the text refer to the same character by analyzing linguistic patterns, contextual cues, and discourse coherence. Do not rely on pre-learned entity associations but rather resolve coreference based on narrative structure."
Personality Understanding
"Based on the character’s dialogue and actions, infer their personality traits. Focus on how they interact with others, their tone, and behavioral tendencies. Avoid using prior memorized descriptions of characters and derive insights purely from the given passage."
Role Detection
"Determine whether a character plays a key role in the narrative (e.g., antagonist, victim, or protagonist) by analyzing their function within the story. Identify how their actions and relationships define their role rather than relying on predefined labels."
Open-Domain QA
"Answer the question using only the given passage. Identify relevant information by reasoning over the text and analyzing interactions between characters. Do not rely on external knowledge or prior memorized facts."
Summarization
"Summarize the passage by identifying key events, character interactions, and narrative progression. Construct a coherent summary that captures the essence of the scene without directly copying text segments."